A Complete Bibliography of Publications in Synthese, 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.03

Title word cross-reference

n [803]. Π_{2}^{1} [460].

-crises [803]. -logic [460]. -terms [384].

10 [599]. 12 [813].

4 [10].

5 [99].

6 [184].

7 [281].

8 [352].

9 [525].

Aarsleff [776]. Abolishing [697]. absent [623]. Absolute [660, 651]. accelerating [179]. acceptance [92, 349]. accessibility [663]. account [103, 659, 904, 332, 639, 567]. Acquaintance [77, 73]. acquisition [224]. action [34, 17, 699, 888, 536, 37]. Actions [569, 258, 530]. Actualism [147]. acuity [169, 170]. additions [907]. Adequacy [157]. adequately [447]. Adjudication [862]. advances [711]. affecting [435]. affirmations [509]. Against [588]. agency [428]. agent [590]. agent-causality [590]. agents [244]. aggregation [451]. agreement [444]. Agreements [277]. Aleatory [135]. algebra [547]. Alien [322]. alleged [279]. almost [268]. alternative [539]. alternatives [803, 292]. amalgamation [301]. ambiguity [380, 279]. analogy [122]. Analysis [816, 729, 587, 466, 474, 905, 589, 801, 405, 140]. analytic [768]. Analyticity [211]. analytics [799]. Anaphora [269, 555]. ancient [923]. animal [498]. Announcement [29, 51, 100, 109, 117, 130, 177, 192, 232, 257, 274, 364, 372, 382, 409, 425, 454, 468, 506, 557, 585, 616, 687, 753, 824, 833, 839, 844, 851, 885, 909, 910, 680, 745, 818, 893]. Announcements [22, 433, 657, 667, 707]. Announcing [453]. anomalism [57]. answered [104]. Anthropology [304, 214]. anti [332, 837, 603]. anti-realism [837]. anti-realist [332]. antimony [123]. antinomies [398, 829]. antirealism [924]. any [68]. Anything [56]. appearances [116]. appendix [829]. Apperception [120]. applicability [789]. application [677, 921]. applications [869]. applied [924]. Applying [556, 925]. approach [737, 395, 878, 757, 869, 9]. approaches [458, 108]. approximate [602, 40]. approximation [438, 63]. apriorism [604]. apt [769]. Archimedean [670]. architecture [871]. Archive [514]. argument [370]. Argumentation [691]. arguments [180, 627]. arising [38]. Aristotelian [279]. Aristotle [278]. Arms [802]. Art [260]. article [20, 219]. artificial [325]. Ascending [617]. ascription [209]. Aspects [472, 323, 470, 615]. assumption [400]. assumptions [613, 189]. asymmetry [806, 589]. Atomic [543]. atomism [76]. atoms [76]. attempt [474]. attitudes [212, 610, 265, 272, 213]. attributing [154]. attributive [153]. Austin [72]. automata [652]. autonomy [717, 343]. available [447]. averaging [691, 347, 446, 451]. Award [641]. awareness [567]. axiomatizing [395]. axioms [670].

back [365]. background [11]. Baker [371]. barefoot [926]. based [681, 103, 661]. bases [23]. basis [227]. Bayesian [349, 346, 594, 89]. be [694, 905, 533, 104, 205, 253, 808, 498, 773]. Beardsley [84, 83]. beauty [835]. became [244]. before [489, 515, 262]. behavior [794, 790, 610, 538]. behaviorism [324]. being [795]. Belief [679, 283, 331, 721, 80, 904, 769, 150]. Beliefs [861]. Believe [355, 353]. believers [583]. believing [860]. Bell [648, 254, 255, 437, 206]. Berkeley [781]. Bertrand [626]. betting [902]. between [815, 593, 495, 739, 780, 550, 356, 603]. Bibliography [917, 524, 884]. binary [400]. Biological [927]. Biologie [522]. Biology [522, 14]. bioprograms [492]. Bivalence [787]. blueprint [871]. blurriness [772]. boat [430]. body [225, 558]. Boolean [547]. Bootstrapping [926]. Born [921]. Botha [163]. bottom [653]. boundaries [712]. breaking [712]. bridge [356]. Bridging [712]. brief [875]. bring [321]. Brown [309]. Bulletin [99, 184, 281, 525, 813, 10, 352, 599]. Burdick [213]. Burley [491]. Burns [897].

calculi [548]. calculus [815, 32, 146, 801, 216]. Call [598, 825, 929]. Can [905, 520, 533, 808, 764, 104, 773]. cannot [267]. can’t [205]. care [241]. careenium [236]. Caring [242]. Carnap [438]. Cartwright [45]. case [708, 143, 254, 400, 63, 924, 477]. Castañeda’s [33]. catalogues [158]. Categories [199, 279, 119]. Causal [676, 727, 806, 609, 589, 183, 613, 333, 32, 146, 611, 588, 808, 482, 639]. Causal-Relevance [183]. Causality [828, 138, 590, 420, 132, 134, 612]. Causation [329, 606, 608, 591, 587, 586, 589, 140, 17]. cause [751]. Causes [874, 141]. century [747]. certain [535]. certainties [810]. Certainty [327, 127]. challenge [782]. Chameleonic [406]. chance [773]. change [622, 908, 642, 470, 615, 883, 907]. charity [634, 723]. Charybdis [437, 206]. chemistry [645, 647, 646]. child [523]. Chinese [325]. Chinese-room [325]. Choice [461, 250, 668, 160, 483, 801]. choices [182]. Chomsky [357]. Chomskyan [230]. Circle [512, 514]. circular [618]. Circularity [625]. Circumstances [482]. City [393]. claims [646]. classical [329, 173, 45]. classificatory [778]. clausally [385]. closures [397]. Coffa [79]. Cognition [881, 716, 569]. Cognitive [664, 928, 856, 671, 768, 663]. Cohen [201]. coherence [619, 629, 863, 792, 293, 494, 522, 522]. coherentism [287]. Collective [888, 721, 668]. college [393]. collision [686]. color [540, 259, 847]. Colour [368]. comes [234]. Comment [576]. Commentary [84]. Comments [169, 375, 248, 92, 246, 377, 53, 243, 94, 126, 239, 235, 166, 237]. commitments [873]. Commodious [701]. Communication [374, 307]. Comparative [734, 836]. comparisons [427]. competency [341]. Complementarity [873, 115]. complemented [385]. complex [771, 337]. complexity [459, 918, 871]. computation [465]. computational [546, 856]. computationalism [654]. computer [108]. concept [542, 862, 190, 9, 563, 458, 516, 854]. conception [560]. conceptions [32, 146, 341, 88]. concepts [622, 322, 785, 784, 822, 368, 12]. Conceptual [722, 572, 195, 883, 590, 622, 895, 521, 912]. conceptualistic [914]. concerning [169]. condition [589]. conditional [648, 553, 253]. conditionalization [477]. Conditionals [907, 478, 791, 253, 548, 773, 787]. Conditions [485, 253, 157]. conditionship [676, 808]. conference [640]. confessions [195]. confirms [56]. conflict [799]. Conflicting [32, 146]. congruence [354]. consciousness [234, 428, 780, 516]. consensual [447, 691]. Consensus [452, 444, 446, 170, 449, 1, 450, 445]. consequence [458]. Conservatism [857]. conservative [774, 397]. considerations [918]. consistency [215]. constitution [888]. Constrained [630]. construction [411, 881]. constructions [113]. Constructive [124, 852]. Constructivism [519]. Contemplation [702]. Contemporary [247]. Content [422, 662, 209, 568, 660, 664, 679]. content-ascription [209]. contents [674, 713]. context [715, 422]. contingently [841]. continuing [799]. continuities [559]. Continuity [602, 507]. contradiction [545]. Controversies [646]. convention [374, 69]. conventions [277]. convergence [922]. Correction [892, 689]. Correlation [586, 176]. correlational [665]. correspondence [62]. Corrigendum [256]. Cosmological [182]. Could [498]. counter [706]. counter-examples [706]. Counterfactual [905, 589, 140]. counterfactuals [794]. Coupling [62]. craft [407]. creativity [357]. crime [535, 926]. criminal [535]. crises [803]. criteria [886, 369]. criterion [675, 215, 126]. critic [580]. Critical [210, 19, 223, 33, 876]. critics [502]. critique [319, 829, 132, 45, 571]. cross [708]. cross-disciplinary [708]. Crusius [779]. current [165, 11]. curriculum [186]. Curse [624]. Cushing [166].

dagger [262]. Darwin [107]. Data [868]. Davidson [375, 894, 895, 572, 637, 37]. dead [577]. debate [921, 705]. debris [705]. Decision [771, 138, 903, 249, 349, 145, 127, 588, 482, 774, 773]. decomposition [251]. deduction [120, 119]. deep [535]. defeasibility [288]. defeasible [554]. defended [145]. Defense [340, 389]. Defenses [774]. definition [636]. deliberation [903]. deliberation-probability [903]. deliberative [792]. democracy [191]. Demonstrative [152, 154, 568]. Dennett [235, 571]. denotation [740]. Denoting [47, 48, 74]. deontic [552, 899, 533, 534, 273]. dependence [288]. derivations [276]. Descartes [580, 779, 780]. descriptions [535, 269]. Destroying [450]. detachment [273]. detection [134]. determination [628]. deterministic [350]. Development [507, 75, 49, 814, 521, 97]. Developmental [42]. d’Everett [180]. Devlin [705]. dia [427]. dia-philosophical [427]. diagnosis [108]. dialectic [829]. Dialectical [102, 103]. Dialectics [4]. dialogical [188]. dialogue [48, 488, 490, 486]. Dialogues [577]. dichotomous [587]. dictionaries [96]. dicto [272]. Did [415, 119]. differences [596, 181]. difficulties [576, 575]. difficulty [345, 346]. Dilemma [302, 635, 614, 793, 481]. direct [224]. directed [538]. directions [486]. disciplinary [708]. discouraged [711]. discourse [543, 741, 269]. discoveries [867]. Discovery [689, 816, 107, 101, 750]. Discussion [163, 240]. disposition [610]. Dispositions [31]. disputation [491]. dissent [801]. distance [593]. distinction [815, 638]. distinctions [654]. diverged [244]. diversity [497, 768]. Division [99, 525, 813, 352, 599]. DL [899]. DLMPS [813, 184, 281, 10]. do [447, 67, 267, 388]. doctrine [535, 320]. Dogmas [186]. Doing [578, 33]. domain [6]. dominance [482]. dose [927]. dose-response [927]. doxastic [864]. Dr [203]. dreams [340]. Dretske [659]. driven [767]. dualism [565]. Dyadic [273, 552]. dynamics [903, 686].

earliest [410]. early [49]. ecological [542, 661]. Ecology [7, 539, 3, 15, 4, 11, 6, 12, 2, 13]. Ecology- [7]. economic [103]. economics [789]. editor [911, 797, 393, 607]. Editorial [52, 403, 418, 426, 455, 282, 289, 185, 164, 172, 178]. education [191, 879, 883, 876]. effects [927]. efficacy [763]. Ehring [485]. Einstein [921, 254, 255]. Einstein-locality [255]. either [841]. element [923]. elementary [872]. Embedding [600]. emergence [709]. Empirical [492, 690, 122, 349, 156, 104, 617]. empirically [234]. Empiricism [604]. empiricist [223]. energy [165]. engineering [925]. entities [562]. entropic [730]. entropy [478, 476, 477]. environment [542]. Epistemic [92, 500, 678, 288, 759, 763, 625]. epistemics [542, 719]. Epistemological [206, 862]. Epistemology [284, 295, 345, 502, 283, 720, 495, 661, 58, 518, 882, 309, 285, 768]. Equilibration [882]. equilibrium [768]. ergodic [251]. Erotetic [748]. Errata [193, 226, 526]. Erratum [28, 499]. error [595]. essay [166, 343]. Essentialism [2]. estimates [927]. Estimation [595, 438, 594]. ethico [682]. ethico-political [682]. ethics [511, 244, 700, 227, 706]. Even [358]. Even-ifs [358]. event [157]. events [143]. Everett [180]. evidence [733, 644, 510]. evidential [677, 903, 774]. evolution [90, 107, 70]. evolutionary [559, 14]. exact [923]. examples [113, 706]. Exclusion [645, 847]. excursion [61]. existence [278, 563, 646]. Existential [559, 560]. expansions [907]. expected [483]. experience [122, 368]. experiential [155]. experiment [325]. experimental [613, 756]. experiments [176]. explain [520]. explanation [34, 921, 540, 473, 142, 472, 143, 538, 823, 183, 144, 152, 136, 628, 134, 137, 42]. explanations [135, 133]. exponent [919]. expressing [860]. expression [190]. extension [179]. Extensive [770, 670]. externalism [761].

facing [582]. Facts [229, 86, 870]. Faithful [670]. fall [376, 710]. fallacy [928, 898]. Fallibilism [619]. father [203]. faveur [180]. favor [180]. features [716]. Feynman [543]. Fiction [82, 414, 836, 84, 83]. fictions [361, 414]. fideists [584]. field [492, 181]. Fifth [685]. figure [628]. figure-ground [628]. Finetti [399]. finite [650]. finitism [197]. Finocchiaro [473]. First [681, 229]. First-order [229]. fit [559]. flaws [863]. flow [442]. Fodor [37]. following [367]. follows [650]. force [731, 690]. form [212, 315, 316]. Formal [488, 630, 470, 615, 490, 207, 451, 97]. Forms [80, 196]. Foundational [465, 456, 467]. Foundationalism [288, 618, 287]. Foundations [719, 405, 326, 462, 645, 156, 604, 617, 306, 671]. Fourth [684]. Fraassen [437]. Frames [431, 179]. Frankfurt [242, 243]. free [595]. Freedom [536]. Frege [427, 816, 376, 398, 896, 826]. French [522, 180, 381, 378]. Friday [489]. Friendship [703]. function [778, 331, 73, 603]. Functionalism [390, 501, 652]. functions [400, 770]. fundamental [654]. Fuzziness [785, 784]. fuzzy [40].

Galileo [580, 564, 474]. gambit [618, 287]. Game [69, 758]. Game-theoretical [758]. games [798, 636, 279, 807]. gap [356]. Gareth [95]. General [677, 789, 869, 601, 873]. generality [159]. generalizability [608]. generalizations [602]. generalized [852]. Generic [744]. Genetic [58, 518]. geometrization [819]. geometry [685]. get [764]. ghosts [244]. Gibb [922]. Gibson [260]. Gideon [480]. Gilbert [631]. give [470, 615]. goal [538]. goal-directed [538]. Gödel [570]. Goodman [86, 53, 87, 54]. grammar [782]. grammars [556]. graph [556]. Gravitation [731]. Greeks [415]. ground [628]. group [483]. growth [107]. guerrilla [800]. Guest [911, 797]. Guise [427].

Habermas [637]. Hacker [371]. half [739]. half-way [739]. Hamann [782]. Hans [776]. Hansson [900]. happiness [702]. happy [711]. Hart [705]. head [861]. Heidegger [817]. Hempel [92]. hidden [44, 730]. high [165]. high-energy [165]. Hintikka [759, 755, 377, 127]. Hintikka-style [759]. historical [642, 583]. History [99, 525, 813, 352, 599, 875, 884, 576, 578, 472, 581, 15, 518, 244, 644, 700, 393, 579, 575]. history-of-science [472]. Hobbes [780]. Hofstadter [237]. holism [420]. homogeneous [890]. hotter [175]. how’s [751]. human [34, 339, 234]. hundred [377, 376]. hunt [926]. hunters [169, 170]. hunters- [170]. Husserl [564, 561, 816, 562, 817, 563]. hydrogeology [924]. Hypertasks [440]. hypotheses [349, 13]. hypothesis [593]. Hypothetical [553].

idea [521]. ideal [452]. idealism [61]. Ideality [534]. idealizability [602]. idealization [564]. Idealized [924, 925]. identical [603]. identification [360]. Identifying [189]. Identity [841, 434, 424, 157, 314, 392, 238]. ideology [379]. ifs [358]. Ignorance [250, 749]. II [344, 139, 172, 162, 504]. III [229, 178]. immediate [286]. impact [88]. imperative [59]. imperatives [553]. implications [179, 206]. implicit [189]. importance [241]. Incognizables [64]. incoherence [897, 620]. incoherent [71]. incompletely [864]. inconsistency [610]. indefinite [729]. independence [301]. Indeterminacy [57, 211]. indeterministic [350]. Indexical [149, 150, 155, 628]. Indexicals [148, 842]. Indian [638]. indicative [773]. indifference [532]. indirect [608]. Indiscernibles [841, 434]. Indiscerniblity [24]. indispensability [762]. indistinguishability [603]. indistinguishables [392]. individual [365]. individualism [611]. individuation [840, 381]. indoctrination [621, 327]. induction [90]. Inductive [137]. ineffable [781]. Inequality [206]. infant [558]. inference [477, 431, 338, 597, 735]. Inferences [361, 484, 530]. infinite [175, 650]. Infinitesimals [783]. influence [376]. informal [898, 461]. Information [661, 663, 447, 662, 677, 103, 659, 398, 595, 666, 136, 679]. Information-based [661]. information-theoretic [659]. informational [660]. inquiring [103]. Inquiry [434, 456, 755, 105, 756, 751, 471, 748, 444]. instantiation [890]. Institutional [337]. institutions [492]. Instrumentalism [91]. intelligence [325]. intensional [737, 744]. intensionality [424]. intent [535]. intention [892]. Intentional [568, 423]. Intentionality [420, 419, 889, 666, 413, 827, 261, 424]. intentions [33, 429]. interdisciplinary [878]. Interest [879, 767]. Interest-driven [767]. Internal [809, 762, 764]. internalist [761]. international [804, 99, 525, 813, 352, 599]. Interpretation [735, 821, 546, 366, 894, 572, 180, 556, 45, 251, 724, 180]. interpretations [125]. Interpreting [821, 383]. interrogative [103, 755, 105]. Interrogatives [471]. Intertheoretic [63]. intertranslatability [845]. interval [593]. Intransitive [532]. intrinsic [434]. Introduction [81, 574, 460, 746, 754, 714, 915, 920, 708]. Introductory [110, 118, 775]. intuitionist [582]. intuitionistic [408]. intuitivity [111]. Investigations [321, 38]. irrational [498]. irrationality [723]. isn’t [194]. Israel [94]. issue [301]. iterated [481]. Iteration [791]. itself [726]. IUHPS [813, 184, 281, 10]. IUHPS/DLMPS [813, 184, 281, 10].

J [260]. J. [677, 789, 260]. James [625]. Jean [522, 520, 523]. John [576]. Jones [899]. journal [453]. judgement [733]. judgment [626]. Justification [293, 297, 717, 496, 762, 834, 291, 618, 448, 765, 68, 294]. justificationist [161]. Justifying [696].

Kant [121, 113, 782, 78, 112, 122, 79, 829, 752, 111, 119, 827, 114, 123, 826]. Kantifying [271]. keep [773]. Keller [625]. Kepler [63]. Keynes [677, 789]. keys [877]. killers [926]. kinds [15]. kinematics [685, 41, 478]. knower [766]. knowing [396]. Knowledge [715, 410, 294, 292, 286, 328, 659, 508, 520, 290, 718, 224, 297, 296, 865, 431, 786, 826, 881]. known [720]. Kochen [44]. Kripke [370, 397, 652, 724].

laboratory [6]. Lacan [558]. Laddaga [301]. Lakatos [641]. lament [679]. Langage [381]. Language [636, 738, 163, 359, 370, 779, 386, 815, 778, 627, 277, 323, 776, 376, 383, 556, 388, 379, 325, 116, 780, 631, 638, 854, 381]. languages [158, 845, 406]. later [376, 377]. Laudan [755]. law [141]. lawlikeness [218]. laws [298, 141, 924, 202, 870]. Learning [581, 190, 668, 877, 359]. lecture [685, 681, 684, 682, 683, 686]. Lectures [543]. Lee [806]. legal [690]. Lehrer [447, 449, 1, 445]. Lehrer/Wagner [447, 449]. Leibniz [779, 778, 776, 777]. Leslie [917]. Leśniewski [398]. lesson [457]. level [927]. levels [618, 127, 287]. Levi [345, 252, 346]. Lewis [140]. lexicon [97]. Liberal [704]. limit [173]. Linguistic [741, 535, 777, 567]. linguistics [640, 711]. lion [895]. lions [895]. list [674, 713]. Literalism [304, 214]. literary [514]. literature [430]. living [701]. local [651, 176]. locality [255]. Locke [776, 777]. logemes [492]. Logic [99, 525, 640, 813, 384, 352, 401, 599, 853, 552, 546, 40, 492, 759, 815, 408, 333, 472, 784, 460, 50, 487, 105, 756, 145, 533, 898, 534, 549, 748, 457, 489, 855, 470, 615, 604, 221, 263, 308, 407, 491, 531, 486, 912, 750]. Logical [38, 76, 108, 550, 535, 789, 212, 805, 49, 816, 315, 814, 918, 31, 556, 316, 190, 458, 519, 204]. logics [492, 734, 228]. look [847]. lost [61]. lottery [200]. low [927].

M [677, 789]. machinery [570]. MacIntyre [245, 246]. made [363, 194]. make [867]. making [154, 249, 127]. man [498]. manifestations [611]. many [821, 549]. many-valued [549]. maps [822]. Margaret [631]. Marginalité. [522]. Marginality [522]. mark [565]. Markovian [612]. marriage [711, 89]. mass [647]. Massey [898]. materialism [389, 573, 30, 2]. Mathematic [913]. Mathematical [457, 789, 439, 332, 853, 786]. mathematician [396]. Mathematics [464, 326, 651, 113, 467, 466, 47, 649, 914]. matter [48]. Matthews [95]. Maximum [477, 476]. Maxwell [573]. may [106]. Mayo [597]. me [321, 262]. Meaning [129, 740, 735, 215, 742, 366, 887, 741, 738, 236, 817, 736, 637, 639, 854, 354, 356, 378]. meaningful [601]. meaningfulness [298, 378]. Meanings [36]. measure [593, 730]. measurement [608, 670, 770, 772]. measures [251]. Measuring [66]. mécanique [180]. mechanical [45]. mechanics [180, 436, 474, 820, 686, 603]. mechanism [805, 151]. medical [127, 108]. meets [873]. membership [815]. memory [339, 297]. Mental [570, 653, 555, 57, 31, 565, 388, 566]. mentalism [23, 230]. Merleau [558, 560]. message [487]. meta [582]. meta-methodologies [582]. Metaknowledge [765]. Metalinguistic [565, 905]. metaphor [360]. metaphors [8, 630]. Metaphysic [915]. metaphysical [650, 195]. Metaphysics [572, 916]. metascience [306]. Metatickles [774]. Method [438, 816, 751, 445, 451]. Methodological [23, 611]. methodologies [165, 582]. Methodology [99, 525, 813, 352, 599, 769, 580, 427]. methods [789, 333]. Michael [633]. microphysical [254, 168]. middle [317]. Mind [662, 225, 887, 234, 244, 655, 247, 501, 359]. Mind-body [225]. minddependent [442]. minds [335]. misrepresentation [880]. missing [430]. mistakes [583]. misunderstanding [370]. Mittgenstein [326]. mixture [7]. modal [229]. Modalities [779, 395, 908]. modality [39]. model [675, 643, 331, 180, 755, 105, 188, 814, 800, 267, 127, 566, 628, 407]. model-theoretic [675, 814]. Modeling [803, 108, 609]. Modelling [721]. Models [790, 165, 218, 176, 642, 927, 926, 870]. modern [88]. Monroe [84]. monsters [36]. Montague [758]. Moral [191, 698, 694, 690, 244, 695, 343]. morality [704, 697, 696]. Moritz [510, 508, 514]. motivation [877]. movements [258]. moves [161]. multi [869]. multi-relation [869]. Multiple [608, 842]. mundane [388]. Must [551]. mutual [904]. mutuality [542]. Myth [913, 914].

names [361]. narrative [86]. National [798]. Natural [15, 539, 383, 556, 533, 845, 201, 541]. naturalism [688]. naturalist [679]. naturalistic [502]. naturalized [661, 205, 285]. Naturalizing [429, 892]. Nature [285, 121, 323, 50, 820, 419, 631, 202]. Near [299]. necessity [201]. negation [545]. Negative [175]. Negativity [918]. Neither [726]. Nelson [86]. nets [556]. Neurath [514, 430]. neuroscience [848]. Newcomb [793, 481, 672]. Newton [755, 63]. niche [542, 8, 9]. No [99, 184, 281, 525, 813, 10, 352, 599, 489, 44, 272]. no-hidden-variable [44]. Non [612, 922, 672, 161]. non-justificationist [161]. Non-Markovian [612]. non-standard [672]. non-uniform [922]. nonexistent [416]. nonindividualistic [566]. nonreferential [404]. nonstandard [466]. nonstatement [68]. nor [726]. Normative [750, 58]. norms [678, 337, 536]. Note [674, 713, 403, 418, 426, 455, 607, 110, 118, 427, 72, 3, 899, 775, 39]. Notes [363]. nothing [234]. notice [210, 19]. notion [49]. notions [341]. Null [13]. number [919]. Numbers [503, 168]. Nyaya [896].

object [681, 457]. Objective [351, 442, 773, 349, 348]. objectivity [350, 597]. objects [416, 744, 258]. obligation [695]. obligations [553, 491]. Oblique [17]. observability [872]. observation [168]. observations [479]. observer [452]. observing [171]. Obvious [328]. Occurrences [217]. occurrents [31]. omissions [531]. omnitemporal [228]. Once [190]. one [808]. one-way [808]. Ongoing [296]. ontological [76, 873]. ontology [675, 24, 853, 260, 536]. opening [321]. operation [919]. opportunity [802]. optimum [438]. order [229, 532]. Ordered [596, 864]. orderings [300, 299, 160]. Organism [542]. Organism-environment [542]. origins [846]. ostensive [636]. other [335, 558]. Otto [514]. Ought [551]. outcome [444]. overview [606].

pairs [308]. Panglossian [501]. paper [84]. papers [598, 825, 929]. Paraconsistent [549]. paradigm [742, 856]. paradigms [3, 2, 61]. paradox [480, 621, 436, 200, 441, 922]. paradox- [480]. Paradoxes [402, 40, 724]. Parfit [239]. parsing [556]. part [344, 172, 178, 131, 336, 139]. partial [586]. particles [872, 777, 603]. particular [143]. particulate [890]. partitives [363]. Parts [363]. past [581, 435]. path [739]. pattern [7]. Pauli [645]. pays [18]. peace [802]. Peacocke [632]. Penelhum [584]. perceived [14]. perceived-randomness [14]. perceiving [261]. Perception [887, 540, 224, 497, 263, 422, 264]. Perceptual [421]. Personal [718, 238]. persons [572]. perspective [712, 346, 318, 504]. perspectives [411]. Peter [230]. Phaedo [874]. phenomena [115, 868]. Phenomenal [392]. Phenomenology [786]. philosophers [880]. Philosophical [321, 642, 341, 30, 877, 75, 427, 878, 819, 524, 179]. Philosophy [99, 525, 813, 352, 599, 576, 49, 578, 466, 581, 829, 244, 875, 884, 73, 435, 247, 638, 309, 501, 393, 914, 647, 575]. Physical [728, 684, 726, 670]. physical/technological [684]. physicalism [159, 390, 513]. physics [329, 467, 165, 159, 819, 435, 543]. Piaget [522, 520, 524, 521, 523]. Piagetian [882]. pictures [334]. pink [890]. place [700, 419]. Plantinga [837]. platonic [48]. Platonism [649]. Plausibility [160, 300, 592]. poem [380]. political [682]. politics [89]. Ponty [558, 560]. Popcorn [903]. population [611]. Pörn [899]. position [789]. possibility [520]. possible [265, 127, 133, 318, 504]. possible-worlds [318]. Practical [345, 346, 923]. practice [681, 684, 682, 613, 853]. Pragmatic [276, 530]. pragmatism [722, 102]. Pre [893]. Pre-announcement [893]. precedent [690]. predator [800]. predicate [384]. predicates [71]. Predication [815]. predictive [12]. Preface [131, 336, 344, 373, 266, 275, 312, 139, 394, 443, 658, 760, 493, 505, 544, 517, 233, 527, 52, 282, 289, 911, 797, 185, 164, 172, 178, 393]. preference [552, 596]. preferences [596]. Present [623]. Preserving [614, 822]. Press [537]. prey [800]. primacy [423]. primitive [590]. Principle [841, 634, 723, 62, 476, 645]. principles [554, 478, 47, 625]. Prior [475, 227]. priori [826]. priority [676]. Prisoner [302, 614, 793, 481]. private [370, 627]. probabilism [399, 2, 7]. Probabilistic [134, 587, 32, 146, 44, 132, 133]. probabilities [447, 249, 349, 902, 728, 253, 45, 475]. Probability [41, 478, 610, 142, 350, 400, 301, 136, 729, 803, 351, 903, 250, 295, 124, 251, 125]. Problem [903, 107, 803, 449, 864, 335, 284, 648, 562, 90, 723, 532, 106, 793, 730, 664, 168, 101, 481, 768]. Problems [508, 672, 467, 879, 471, 582, 465, 752, 11, 706]. procedure [691]. processes [564]. processing [677]. Professor [87]. profile [155]. program [609, 819]. programming [38]. Programs [386]. Progress [97, 65]. progressive [23]. Prolegomena [535]. Proof [459, 463, 44, 733, 405, 649]. propensities [143]. propensity [804]. properties [434, 445, 623, 451]. proposal [1]. proposition [733]. propositional [212]. Propositions [59, 807]. Prototypes [784]. Provability [439]. pseudo [217]. pseudo-occurrences [217]. psychiatric [341]. psycho [511]. psycho-sociological [511]. Psycholinguistics [712, 708, 709, 710]. Psychological [886, 152, 163]. psychologism [656, 58]. psychologist [523]. psychology [856, 495, 518, 711, 671, 343, 338]. psychophysiological [540]. psychosemantics [665]. PTQ [38]. Publications [16]. publisher [674, 713, 745, 818]. Pufendorf [700]. puzzle [355, 354].

Qualia [540, 624]. qualitative [181]. qualities [392]. quantification [383]. Quantified [759]. quantifiers [395, 413, 404, 852]. quantique [180]. Quantitative [181, 298]. quantum [173, 180, 436, 820, 604, 45, 168, 308, 603]. Quark [168, 169, 170]. quarks [171]. quasi [613]. quasi-experimental [613]. Quelques [180]. question [323]. questioning [747]. questions [758, 187, 757, 104]. Quine [635, 330, 430]. quotients [902].

races [802]. radiation [927]. radical [380]. ramification [180]. Ramsey [791]. random [331, 479, 795]. Randomness [14, 254]. range [536]. ranking [592]. rapid [710]. ratificationism [774]. ratio [770]. Rational [691, 749, 473, 250, 445]. Rationality [339, 347, 695, 483, 343, 338, 528, 345, 480, 643, 342, 346, 341, 340, 238, 88, 698, 876, 337, 70]. Rationalization [528, 637]. reaching [445]. reading [72]. ready [194]. ready-made [194]. Real [290]. Realism [906, 208, 725, 207, 222, 221, 220, 619, 225, 223, 809, 122, 196, 739, 61, 837, 206]. Realist [209, 195, 332, 260]. Reality [316, 779, 163]. Realizability [408]. really [415]. Reason [694, 782, 835, 205]. reasonableness [342]. Reasoning [554, 855, 40, 690, 401, 767, 519]. reasons [17, 496]. received [16]. reconsideration [702]. reconsidered [144]. reconstruction [511]. Red [843, 632]. reduction [822]. Reductionism [541, 4]. Reductive [125]. Reference [414, 154, 866, 155, 151, 421, 222, 628, 93, 148, 638, 149, 356]. Referential [215, 404, 153]. Referential/ [153]. Reflections [581, 792, 123, 430]. Reflective [908, 768]. Reformulation [730]. regained [61]. regarding [674, 713]. regulating [720]. Reichenbach [183]. rejoinder [900, 355]. relata [727, 806]. Relation [751, 677, 495, 183, 869, 780, 603]. Relational [407, 600]. relations [550, 125]. relative [384]. Relativism [643, 688, 620]. relativity [873, 179]. release [537]. Relevance [183, 303]. Relevant [743, 466, 548, 292]. reliability [294, 496]. remains [514]. Remarks [456, 321, 509, 644, 458, 326, 816, 315, 819, 736]. replies [502]. Reply [245, 806, 252, 633, 95, 5, 485, 242, 230, 487, 898, 301, 897, 625]. Representation [317, 561, 413, 579, 82, 741, 653, 601, 670, 84, 83]. representational [904, 359]. representations [269]. representing [864]. requirements [38]. research [708, 790, 642]. resolution [40, 398]. resonance [646]. respect [446]. Response [310, 83, 86, 167, 927, 170, 597]. responses [354]. results [552]. retheoretization [62]. reticulation [688]. reticulational [643]. retrieval [663]. reversal [307]. Review [849, 432, 812, 280, 26, 305, 858, 35, 901, 417, 198, 859, 27, 832, 25, 231, 838, 60, 796, 788, 831, 20, 891, 219, 673, 311, 811, 732, 21, 605, 830, 850, 875]. Revisions [907, 774]. revisited [325]. revitalized [905]. revival [747]. revolution [98]. rich [773]. right [268, 669, 696]. rigour [461]. rise [376, 710]. risk [249, 927]. Ritual [93]. role [581]. room [325]. Rorty [225, 561, 248]. Routines [529]. Rudolf [163]. rule [367]. Rules [113, 366, 319, 488]. Russell [268, 75, 49, 48, 78, 50, 79, 46, 410, 411, 73, 626].

sacred [98]. Sailing [437]. Salmon [183]. sameness [181]. Saturated [896]. saying [860]. Scales [298]. Sceptics [583]. Scheffler [94]. Schemes [572, 895, 878]. scheming [195]. Schlick [223, 507, 508, 511, 514, 515, 865, 510]. scholium [474]. school [191]. Science [99, 525, 813, 352, 599, 681, 866, 880, 789, 67, 539, 928, 879, 333, 472, 581, 878, 518, 829, 457, 875, 884, 883, 61, 88, 836, 560, 882, 664, 924, 309, 876, 877, 89, 337, 393, 204]. sciences [608, 606, 923, 612, 870, 97]. Scientific [642, 65, 101, 580, 345, 834, 643, 755, 105, 346, 471, 748, 925, 91, 835]. search [750]. searle [386, 889]. Second [682, 369, 122, 123]. sections [321]. security [798]. see [262]. seeing [412]. seeing-a [412]. selection [834, 541]. Self [428, 558, 569, 510]. Self-consciousness [428]. self-evidence [510]. Sellars [622, 567]. Semantic [630, 548, 474, 556, 664, 216, 679]. Semantical [279]. Semantics [385, 856, 387, 737, 758, 398, 128, 267, 209, 207, 264, 318, 504]. semi [532]. semi-order [532]. semiotic [155]. sensation [412]. sensationism [112]. sense [638, 149]. sense-reference [638]. sentences [263, 861, 37]. sentience [692]. separate [921]. sequences [461]. serial [926]. Series [393]. seriously [469]. set [104, 463]. Sets [162, 199, 400, 592]. sets- [400]. shapes [258]. shared [444]. should [187]. shoves [236]. sides [877]. Siegel [310]. signifiance [378]. significance [251]. Signification [378]. Simberloff [3]. Simplicity [872]. Single [143]. Singular [611]. situation [264]. Situations [265, 743, 350]. Sixth [686]. skeptical [502]. Skepticism [291, 91]. skeptics [584]. slab [321]. Slezak [230]. smart [773]. Smith [633]. smoking [773]. Sobel [903]. Social [668, 683, 608, 789, 715, 333, 720, 323, 719, 606, 716, 718, 160, 631, 612, 97]. sociality [717]. sociological [511]. solution [803, 621, 200]. solve [106]. solving [284, 107, 101]. Some [552, 467, 705, 176, 323, 816, 315, 918, 228, 479, 819, 716, 582, 445, 736, 430, 706, 724, 654, 470, 615, 11, 458, 854, 612, 179, 180, 576, 575]. Sorites [362, 441]. space [411, 600, 174, 318, 504]. spaces [299]. special [477, 179]. specific [334]. Specker [44]. specular [558]. Spin [174]. Spinoza [335, 650]. Stalnaker [648]. Standard [672]. standards [715]. starmaking [55]. statements [729, 141, 592]. states [31]. static [866]. stationary [251]. Statistical [144, 789, 348, 484, 597]. statistics [683]. status [484, 114, 476]. step [840]. Stich [671]. Stochastic [255]. stories [87]. story [85, 711]. Strategic [794, 738]. Strengthening [254]. Strict [197]. strikes [365]. strong [228]. Structuralist [822]. structure [535, 692, 339, 129, 549, 822, 490]. structure-preserving [822]. structures [587]. studies [393]. study [234, 85, 33, 836, 873, 163, 204]. stuff [669]. stupidity [763]. style [759, 387]. sub [534]. sub-ideality [534]. subject [49]. Subjective [902, 349]. Subjectivity [313, 597]. subjunctive [787]. substances [647]. substitutional [46, 404]. success [464]. succession [3, 2]. suggested [467]. suggestion [576]. suggestions [575]. summarize [447]. Summary [240]. Supervenience [823, 159, 888]. Suppes [355, 132]. surprisingly [710]. survey [790, 470, 615]. Survival [391]. suspension [810]. syllogism [547]. Symbolism [304, 214]. symmetric [395]. symmetry [603]. SYMOs [870]. symphony [85]. Symposium [708]. Synonymy [566]. syntactical [39]. syntax [535]. System [752, 103, 650, 899]. System-problems [752]. systemic [870]. Systems [9, 611, 864, 488, 869, 308, 407, 804, 750].

Taking [469, 249]. tales [85]. talk [895]. Talking [895]. targets [801]. teach [187]. teachers [880]. teaching [188, 878, 875, 884, 882, 881]. Technical [863]. technological [684]. Technology [928, 683, 393]. Teleology [805]. temperatures [175]. temporary [810]. terms [384, 886, 592, 647]. terrorist [801]. test [791]. testing [348]. tests [869].

TETRAD [609]. Tharp [912, 917, 913]. their [611, 400]. Themes [878]. themselves [828, 116, 114, 115]. Theorem [437]. theorems [255, 612, 916]. theoretic [675, 659, 814, 267]. theoretical [165, 758, 562, 11, 926, 68]. theoreticity [43, 126]. Theories [158, 915, 870, 684, 469, 834, 229, 330, 107, 442, 925, 600, 68, 736, 261, 550, 774, 854]. theorizing [330]. Theory [410, 681, 682, 866, 677, 173, 113, 427, 459, 906, 49, 103, 863, 351, 903, 591, 92, 449, 908, 69, 284, 741, 905, 349, 376, 814, 46, 630, 183, 465, 771, 470, 615, 104, 348, 224, 181, 601, 670, 132, 588, 907, 873, 148, 865, 45, 405, 482, 463, 646, 773, 626, 359]. therapy [225]. There [272, 576, 416, 739, 533, 448, 68, 194, 808, 302, 579, 575]. thesis [889]. things [828, 363, 388, 116, 114, 115]. Thinking [33, 572, 876]. Third [683]. Thisness [270, 147]. thought [779, 507, 325, 152, 568]. thoughts [369]. Three [196, 915, 916]. time [442, 307]. Token [159, 390]. Tomberlin [427]. Tool [457]. topological [299]. toposes [199]. Tractatus [317, 319, 313]. Traditional [502]. transcendental [120, 119]. Transference [630, 485, 591]. transition [45]. transitivity [612]. translation [723, 211]. Translational [128]. treatise [491]. treatments [39]. tribute [693]. true [841]. Truth [733, 835, 649, 545, 546, 67, 384, 439, 593, 293, 18, 332, 548, 620]. Truthlikeness [594, 66]. Tuomela [34]. Turns [90]. twentieth [747]. Twisted [85]. Two [8, 186, 261, 877, 40, 119, 161]. type [389]. types [49, 807].

u [448]. unavoidably [334]. uncertainties [771]. uncertainty [730]. Undefeated [765]. underdetermination [906, 57]. understanding [386, 129, 738, 556, 637, 220]. unexploitable [793]. unification [873]. unified [181]. uniform [922, 726]. uniformity [121]. unimportance [795]. Union [99, 525, 813, 352, 599]. uniqueness [600]. unity [516]. universal [731]. universals [224]. unmirrored [285]. Unreliable [249]. unsaturated [896]. unswept [705]. upon [103]. use [644, 320]. Useful [12]. uses [842]. using [40, 730]. utility [138, 331, 92, 483].

vague [71]. Vagueness [629, 897, 772, 270]. validity [733]. value [500]. valued [549]. Values [927, 216]. variable [44, 730]. Varieties [656]. Verbs [699, 385]. Verificationism [639, 739, 846]. verificationistic [736]. verisimilitude [303, 201]. version [44]. versions [119, 481]. versus [712, 485, 815, 442]. Vienna [512, 514]. view [319, 68]. viewpoint [814]. virtue [494]. virtues [857]. Visual [318, 504, 169, 692, 170]. Viviani [474]. volume [674, 713]. voting [790]. vs [561, 928, 926].

Wagner [447, 449, 445]. Walter [491]. want [321]. war [803, 804]. warfare [800]. was [268, 203]. wave [603]. way [720, 739, 156, 808]. Ways [53, 696]. weight [677, 449]. weighted [691, 347, 446, 451]. Where [921, 416]. whether [323]. while [926]. Who [236, 203, 203]. whom [236]. why-questions [757]. winch [323]. Wisconsin [387]. without [596].

Wittgenstein [317, 370, 366, 371, 412, 314, 627, 323, 319, 410, 848, 324, 367, 512, 325, 515, 847, 320, 724, 846]. Wittgensteinian [315]. wonder [877]. wonderful [54]. word [236]. words [388]. work [524]. Working [462]. world [726, 61, 194, 504]. Worldmaking [53]. worlds [821, 265, 127, 836, 54, 318]. worship [415]. Would [353]. Wright [803]. write [119]. writings [917, 317, 474, 410]. wrong [286, 665, 696].

years [377, 376]. Yolton [576].

zero [449]. Zeus [415].

References

Laddaga:1977:LCP

[1] Robert Laddaga. Lehrer and the consensus proposal. Synthese, 36(4): 473–477, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486109. See reply [301].

Simberloff:1980:SPE

[2] Daniel Simberloff. A succession of paradigms in ecology: Essentialism to materialism and probabilism. Synthese, 43(1):3–39, January 1980. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413854.

Grene:1980:NSS

[3] Marjorie Grene. A note on Simberloff’s ‘succession of paradigms in ecology’. Synthese, 43(1):41–45, January 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413855.

Levins:1980:DRE

[4] Richard Levins and Richard Lewontin. Dialectics and reductionism in ecology. Synthese, 43(1):47–78, January 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413856.

Simberloff:1980:R

[5] Daniel Simberloff. Reply. Synthese, 43(1):79–93, January 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413857.

Mertz:1980:DLE

[6] David B. Mertz and David E. McCauley. The domain of laboratory ecology. Synthese, 43(1):95–110, January 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413858.

Southwood:1980:EMP

[7] T. R. E. Southwood. Ecology- a mixture of pattern and probabilism. Synthese, 43(1):111–122, January 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413859.

Haefner:1980:TMN

[8] James W. Haefner. Two metaphors of the niche. Synthese, 43(1):123– 153, January 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413860.

Patten:1980:SAC

[9] B. C. Patten and G. T. Auble. Systems approach to the concept of niche. Synthese, 43(1):155–181, January 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413861.

Fenstad:1980:IDB

[10] Jens Erik Fenstad. IUHPS/DLMPS Bulletin No. 4. Synthese, 43(1):183– 192, January 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413862.

McIntosh:1980:BSC

[11] Robert P. McIntosh. The background and some current problems of theoretical ecology. Synthese, 43(2):195–255, February 1980. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413926.

Peters:1980:UCP

[12] Robert Henry Peters. Useful concepts for predictive ecology. Synthese, 43 (2):257–269, February 1980. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413927.

Strong:1980:NHE

[13] Donald R. Strong, Jr. Null hypotheses in ecology. Synthese, 43(2):271– 285, February 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413928.

Wimsatt:1980:RPR

[14] William C. Wimsatt. Randomness and perceived-randomness in evolutionary biology. Synthese, 43(2):287–329, February 1980. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413929.

Kiester:1980:NKN

[15] A. Ross Kiester. Natural kinds, natural history and ecology. Synthese, 43 (2):331–342, February 1980. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413930.

Anonymous:1980:PR

[16] Anonymous. Publications received. Synthese, 43(2):343–349, February 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413931.

Stoutland:1980:OCR

[17] Frederick Stoutland. Oblique causation and reasons for action. Synthese, 43(3):351–367, March 1980. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00870293.

Loewer:1980:TP

[18] Barry Loewer. The truth pays. Synthese, 43(3):369–380, March 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00870294.

Miller:1980:CN

[19] David Miller, Catherine Z. Elgin, Jonathan E. Adler, and Douglas N. Walton. Critical notice. Synthese, 43(3):381–431, March 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00870295.

Nelson:1980:RA

[20] R. J. Nelson. Review article. Synthese, 43(3):433–451, March 1980. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00870296.

Sosa:1980:R

[21] Ernest Sosa and Richard E. Grandy. Review. Synthese, 43(3):453– 464, March 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00870297.

Anonymous:1980:Aa

[22] Anonymous. Announcements. Synthese, 43(3):465–468, March 1980. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00870298.

Botha:1980:MBP

[23] Rudolf P. Botha. Methodological bases of a progressive mentalism. Synthese, 44(1):1–112, May 1980. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.

1007/BF00413715. See comment [163].

Kraut:1980:IO

[24] Robert Kraut. Indiscerniblity and ontology. Synthese, 44(1):113–135, May 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413716.

Imbert:1980:R

[25] Claude Imbert and Marlena G. Corcoran. Review. Synthese, 44(1): 137–147, May 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413717.

Blackburn:1980:R

[26] Simon Blackburn. Review. Synthese, 44(1):149–159, May 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413718.

Gardenfors:1980:R

[27] Peter Gärdenfors. Review. Synthese, 44(1):161–169, May 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413719.

Anonymous:1980:E

[28] Anonymous. Erratum. Synthese, 44(1):171, May 1980. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413720;   http://link.
springer.com/content/pdf/10.1007/BF00413720.pdf.

Anonymous:1980:Ab

[29] Anonymous. Announcement. Synthese, 44(1):172, May 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413721.

McGinn:1980:PM

[30] Colin McGinn. Philosophical materialism. Synthese, 44(2):173–206, June 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413406.

Hoy:1980:DLS

[31] Ronald C. Hoy. Dispositions, logical states, and mental occurrents. Synthese, 44(2):207–239, June 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413407.

Fetzer:1980:PCC

[32] James H. Fetzer and Donald E. Nute. A probabilistic causal calculus: Conflicting conceptions. Synthese, 44(2):241–246, June 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413408.

Gustafson:1980:CNI

[33] Donald Gustafson. Castañeda’s intentions: A critical study of Castañeda’s thinking and doing. Synthese, 44(2):247–284, June 1980. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413409.

Audi:1980:TEH

[34] Robert Audi. Tuomela on the explanation of human action. Synthese, 44(2):285–306, June 1980. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413410.

Campbell:1980:R

[35] Keith Campbell. Review. Synthese, 44(2):307–311, June 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413411.

Margalit:1980:MM

[36] Avishai Margalit. Meanings and monsters. Synthese, 44(3):313–346, July 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413465.

Wierenga:1980:FDA

[37] Edward Wierenga. Fodor on Davidson on action sentences. Synthese, 44(3):347–359, July 1980. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413466.

Janssen:1980:LIP

[38] Theo M. V. Janssen. Logical investigations on PTQ arising from programming requirements. Synthese, 44(3):361–390, July 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413467.

Thomason:1980:NST

[39] Richmond H. Thomason. A note on syntactical treatments of modality. Synthese, 44(3):391–395, July 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413468.

Baldwin:1980:RTP

[40] J. F. Baldwin and N. C. F. Guild. The resolution of two paradoxes by approximate reasoning using a fuzzy logic. Synthese, 44(3):397–420, July 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413469.

Domotor:1980:PK

[41] Zoltan Domotor, Mario Zanotti, and Henson Graves. Probability kinematics. Synthese, 44(3):421–442, July 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413470.

Woodward:1980:DE

[42] James Woodward. Developmental explanation. Synthese, 44(3):443– 466, July 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413471.

Balzer:1980:T

[43] Wolfgang Balzer and C. Ulises Moulines. On theoreticity. Synthese, 44(3): 467–494, July 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413472.

Hellman:1980:PVK

[44] Geoffrey Hellman. A probabilistic version of the Kochen–Specker nohidden-variable proof. Synthese, 44(3):495–500, July 1980. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413473.

Shiva:1980:QMT

[45] Vandana Shiva. Are quantum mechanical transition probabilities classical? A critique of Cartwright’s interpretation of quantum theory. Synthese, 44(3):501–508, July 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413474.

Hylton:1980:RST

[46] Peter Hylton. Russell’s substitutional theory. Synthese, 45(1):1–31, September 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413984.

Hursthouse:1980:DPM

[47] Rosalind Hursthouse. Denoting in the principles of mathematics. Synthese, 45(1):33–42, September 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413985.

Coffa:1980:RPD

[48] J. Alberto Coffa. Russell as a platonic dialogue: The matter of denoting. Synthese, 45(1):43–70, September 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413986.

Cocchiarella:1980:DTL

[49] Nino Cocchiarella. The development of the theory of logical types and the notion of a logical subject in Russell’s early philosophy. Synthese, 45 (1):71–115, September 1980. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413987.

Griffin:1980:RNL

[50] Nicholas Griffin. Russell on the nature of logic (1903–1913). Synthese, 45 (1):117–188, September 1980. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413988.

Anonymous:1980:Ac

[51] Anonymous. Announcement. Synthese, 45(1):189, September 1980. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413989.

Anonymous:1980:EP

[52] Anonymous. Editorial preface. Synthese, 45(2):191, October 1980. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413557.

Hempel:1980:CGW

[53] Carl G. Hempel. Comments on Goodman’s ways of worldmaking. Synthese, 45(2):193–199, October 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413558.

Scheffler:1980:WWG

[54] Israel Scheffler. The wonderful worlds of Goodman. Synthese, 45(2):201– 209, October 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413559.

Goodman:1980:S

[55] Nelson Goodman. On starmaking. Synthese, 45(2):211–215, October 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413560.

Hendry:1980:ACA

[56] Herbert E. Hendry and James E. Roper. Anything confirms anything? Synthese, 45(2):217–232, October 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413561.

Elgin:1980:IUA

[57] C. Z. Elgin. Indeterminacy, underdetermination, and the anomalism of the mental. Synthese, 45(2):233–255, October 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413562.

Kitchener:1980:GEN

[58] Richard F. Kitchener. Genetic epistemology, normative epistemology, and psychologism. Synthese, 45(2):257–280, October 1980. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413563.

Huntley:1980:PI

[59] Martin Huntley. Propositions and the imperative. Synthese, 45(2):281– 310, October 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413564.

Korner:1980:R

[60] S. Körner. Review. Synthese, 45(2):311–315, October 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413565.

Nola:1980:PLW

[61] Robert Nola. ‘Paradigms lost, or the world regained’ — an excursion into realism and idealism in science. Synthese, 45(3):317–350, November 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02221783.

Rosenberg:1980:CRC

[62] Jay F. Rosenberg. Coupling, retheoretization, and the correspondence principle. Synthese, 45(3):351–385, November 1980. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF02221784.

Moulines:1980:IAK

[63] C. Ulises Moulines. Intertheoretic approximation: the Kepler–Newton case. Synthese, 45(3):387–412, November 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF02221785.

Levi:1980:I

[64] Isaac Levi. Incognizables. Synthese, 45(3):413–426, November 1980. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF02221786.

Niiniluoto:1980:SP

[65] Ilkka Niiniluoto. Scientific progress. Synthese, 45(3):427–462, November 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF02221787.

Rosenkrantz:1980:MT

[66] R. D. Rosenkrantz. Measuring truthlikeness. Synthese, 45(3):463–487, November 1980. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF02221788.

Cohen:1980:WSD

[67] L. Jonathan Cohen. What has science to do with truth? Synthese, 45 (3):489–498, November 1980. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF02221789.

Pearce:1981:TTJ

[68] David Pearce. Is there any theoretical justification for a nonstatement view of theories? Synthese, 46(1):1–39, January 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064465.

Gilbert:1981:GTC

[69] Margaret Gilbert. Game theory and convention. Synthese, 46(1):41– 93, January 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064466.

Sober:1981:ER

[70] Elliott Sober. The evolution of rationality. Synthese, 46(1):95–120, January 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064467.

Peacocke:1981:VPI

[71] Christopher Peacocke. Are vague predicates incoherent? Synthese, 46 (1):121–141, January 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064468.

Graham:1981:NRA

[72] Keith Graham. A note on reading Austin. Synthese, 46(1):143–147, January 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064469.

Pears:1981:FAR

[73] David Pears. The function of acquaintance in Russell’s philosophy. Synthese, 46(2):149–166, February 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064385.

Hintikka:1981:DW

[74] Jaakko Hintikka. On denoting what? Synthese, 46(2):167–183, February 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064386.

Cappio:1981:RPD

[75] James Cappio. Russell’s philosophical development. Synthese, 46(2):185– 205, February 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064387.

Lycan:1981:LAO

[76] William Lycan. Logical atomism and ontological atoms. Synthese, 46 (2):207–229, February 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064388.

Clark:1981:A

[77] Romane Clark. Acquaintance. Synthese, 46(2):231–246, February 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064389.

Coffa:1981:RK

[78] J. Alberto Coffa. Russell and Kant. Synthese, 46(2):247–263, February 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064390.

Hintikka:1981:RKC

[79] Jaakko Hintikka. Russell, Kant, and Coffa. Synthese, 46(2):265–270, February 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064391.

Grandy:1981:FB

[80] Richard E. Grandy. Forms of belief. Synthese, 46(2):271–284, February 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064392.

Barrett:1981:I

[81] Robert B. Barrett. Introduction. Synthese, 46(3):287–290, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01130041.

Beardsley:1981:FR

[82] Monroe C. Beardsley. Fiction as representation. Synthese, 46(3):291– 313, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF01130042. See comments [84, 83].

Wolterstorff:1981:RBF

[83] Nicholas Wolterstorff. Response to Beardsley on ‘Fiction as representation’. Synthese, 46(3):315–323, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01130043. See [82].

Krukowski:1981:CMB

[84] Lucian Krukowski. Commentary on Monroe Beardsley’s paper, ‘Fiction as representation’. Synthese, 46(3):325–330, March 1981. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01130044. See [82].

Goodman:1981:TTS

[85] Nelson Goodman. Twisted tales; or story, study, and symphony. Synthese, 46(3):331–349, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01130045.

Cohen:1981:FNR

[86] Ted Cohen. The facts of narrative: A response to Nelson Goodman. Synthese, 46(3):351–354, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01130046.

Moreno:1981:PGS

[87] Jonathan D. Moreno. Professor Goodman’s stories. Synthese, 46(3):355– 358, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01130047.

Putnam:1981:ISM

[88] Hilary Putnam. The impact of science on modern conceptions of rationality. Synthese, 46(3):359–382, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01130048.

Simon:1981:BMS

[89] Thomas W. Simon. A Bayesian marriage of science and politics. Synthese, 46(3):383–387, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01130049.

Hempel:1981:TEP

[90] Carl G. Hempel. Turns in the evolution of the problem of induction. Synthese, 46(3):389–404, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01130050.

Leach:1981:ISS

[91] Jim Leach. Instrumentalism and scientific skepticism. Synthese, 46(3): 405–412, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01130051.

Feleppa:1981:EUT

[92] Robert Feleppa. Epistemic utility and theory acceptance: Comments on Hempel. Synthese, 46(3):413–420, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01130052.

Scheffler:1981:RR

[93] Israel Scheffler. Ritual and reference. Synthese, 46(3):421–437, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01130053.

Matthews:1981:CIS

[94] Gareth B. Matthews. Comments on Israel Scheffler. Synthese, 46(3): 439–444, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF01130054. See [93] and reply [95].

Scheffler:1981:RGM

[95] Israel Scheffler. Reply to Gareth Matthews. Synthese, 46(3):445– 448, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF01130055. See [93, 94].

Churchman:1981:D

[96] C. West Churchman. On dictionaries. Synthese, 46(3):449–454, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01130056.

Wolfson:1981:PDF

[97] Robert J. Wolfson. Progress in the development of a formal lexicon for the social sciences. Synthese, 46(3):455–465, March 1981. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01130057.

Marchal:1981:SR

[98] Joe Marchal. The sacred revolution. Synthese, 46(3):467–482, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01130058.

Anonymous:1981:IUH

[99] Anonymous. International Union of the History and Philosophy of Science Division of Logic, Methodology and Philosophy of Science Bulletin No. 5. Synthese, 46(3):483–484, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01130059.

Anonymous:1981:Aa

[100] Anonymous. Announcement. Synthese, 46(3):485, March 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01130060.

Simon:1981:SDP

[101] Herbert A. Simon, Patrick W. Langley, and Gary L. Bradshaw. Scientific discovery as problem solving. Synthese, 47(1):1–27, April 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064262.

Mitroff:1981:DP

[102] Ian I. Mitroff and Richard O. Mason. Dialectical pragmatism. Synthese, 47(1):29–42, April 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064263.

Dacey:1981:IAD

[103] Raymond Dacey. An interrogative account of the dialectical inquiring system based upon the economic theory of information. Synthese, 47(1): 43–55, April 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064264.

Materna:1981:SEQ

[104] Pavel Materna. The set of empirical questions that can be answered by an empirical theory. Synthese, 47(1):57–68, April 1981. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064265.

Hintikka:1981:LIM

[105] Jaakko Hintikka. On the logic of an interrogative model of scientific inquiry. Synthese, 47(1):69–83, April 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064266.

Nickles:1981:WPW

[106] Thomas Nickles. What is a problem that we may solve it? Synthese, 47(1):85–118, April 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064267.

Kleiner:1981:PSD

[107] Scott A. Kleiner. Problem solving and discovery in the growth of Darwin’s theories of evolution. Synthese, 47(1):119–162, April 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064268.

Schaffner:1981:MMD

[108] Kenneth F. Schaffner. Modeling medical diagnosis: Logical and computer approaches. Synthese, 47(1):163–199, April 1981. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064269.

Anonymous:1981:Ab

[109] Anonymous. Announcement. Synthese, 47(1):i, April 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064270.

Butts:1981:INa

[110] Robert E. Butts. Introductory note. Synthese, 47(2):201, May 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064344.

Meerbote:1981:KI

[111] Ralf Meerbote. Kant on intuitivity. Synthese, 47(2):203–228, May 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064345.

George:1981:KS

[112] Rolf George. Kant’s sensationism. Synthese, 47(2):229–255, May 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064346.

Butts:1981:REC

[113] Robert E. Butts. Rules, examples and constructions Kant’s theory of mathematics. Synthese, 47(2):257–288, May 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064347.

Rescher:1981:STT

[114] Nicholas Rescher. On the status of “things in themselves” in Kant. Synthese, 47(2):289–299, May 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064348.

Werkmeister:1981:CPT

[115] W. H. Werkmeister. The complementarity of phenomena and things in themselves. Synthese, 47(2):301–311, May 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064349.

Posy:1981:LAT

[116] Carl J. Posy. The language of appearances and things in themselves. Synthese, 47(2):313–352, May 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064350.

Anonymous:1981:Ac

[117] Anonymous. Announcement. Synthese, 47(2):353, May 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064351.

Butts:1981:INb

[118] Robert E. Butts. Introductory note. Synthese, 47(3):355, June 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01073775.

Meyer:1981:WDK

[119] Michel Meyer. Why did Kant write two versions of the transcendental deduction of the categories? Synthese, 47(3):357–383, June 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01073776.

Howell:1981:ATD

[120] Robert Howell. Apperception and the 1787 transcendental deduction. Synthese, 47(3):385–448, June 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01073777.

Beck:1981:KUN

[121] Lewis White Beck. Kant on the uniformity of nature. Synthese, 47(3): 449–464, June 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01073778.

Harper:1981:KER

[122] William Harper. Kant’s empirical realism and the second analogy of experience. Synthese, 47(3):465–480, June 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01073779.

VanCleve:1981:RKS

[123] James Van Cleve. Reflections on Kant’s second antimony. Synthese, 47 (3):481–494, June 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01073780.

Shafer:1981:CP

[124] Glenn Shafer. Constructive probability. Synthese, 48(1):1–60, July 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01064627.

vonPlato:1981:RRI

[125] Jan von Plato. Reductive relations in interpretations of probability. Synthese, 48(1):61–75, July 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064628.

Pearce:1981:CCT

[126] David Pearce. Comments on a criterion of theoreticity. Synthese, 48 (1):77–86, July 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064629.

Moore:1981:HPW

[127] G. William Moore and Grover M. Hutchins. A Hintikka possible worlds model for certainty levels in medical decision making. Synthese, 48(1): 87–119, July 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064630.

Lepore:1981:TS

[128] Ernest Lepore and Barry Loewer. Translational semantics. Synthese, 48(1):121–133, July 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064631.

Davies:1981:MSU

[129] Martin Davies. Meaning, structure and understanding. Synthese, 48(1): 135–161, July 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064632.

Anonymous:1981:Ad

[130] Anonymous. Announcement. Synthese, 48(1):163, July 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064633.

F:1981:PP

[131] J. H. F. Preface to Part I. Synthese, 48(2):166, August 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063884.

Otte:1981:CST

[132] Richard Otte. A critique of Suppes’ theory of probabilistic causality. Synthese, 48(2):167–189, August 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063885.

Suppes:1981:WPE

[133] Patrick Suppes and Mario Zanotti. When are probabilistic explanations possible? Synthese, 48(2):191–199, August 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01063886.

Rogers:1981:PCE

[134] Ben Rogers. Probabilistic causality, explanation, and detection. Synthese, 48(2):201–223, August 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063887.

Humphreys:1981:AE

[135] Paul Humphreys. Aleatory explanations. Synthese, 48(2):225–232, August 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01063888.

Railton:1981:PEI

[136] Peter Railton. Probability, explanation, and information. Synthese, 48 (2):233–256, August 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063889.

Tuomela:1981:IE

[137] Raimo Tuomela. Inductive explanation. Synthese, 48(2):257–294, August 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01063890.

Eells:1981:CUD

[138] Ellery Eells. Causality, utility, and decision. Synthese, 48(2):295– 329, August 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063891.

JHF:1981:PPI

[139] JHF. Preface to Part II. Synthese, 48(3):332, September 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063983.

Stern:1981:LCA

[140] Cindy D. Stern. Lewis’ counterfactual analysis of causation. Synthese, 48 (3):333–345, September 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063984.

Nute:1981:CLL

[141] Donald Nute. Causes, laws, and law statements. Synthese, 48(3):347– 369, September 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063985.

Fetzer:1981:PE

[142] James H. Fetzer. Probability and explanation. Synthese, 48(3):371–408, September 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063986.

Hanna:1981:SCP

[143] Joseph F. Hanna. Single case propensities and the explanation of particular events. Synthese, 48(3):409–436, September 1981. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063987.

Niiniluoto:1981:SER

[144] Ilkka Niiniluoto. Statistical explanation reconsidered. Synthese, 48(3): 437–472, September 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063988.

Jeffrey:1981:LDD

[145] Richard Jeffrey. The logic of decision defended. Synthese, 48(3):473– 492, September 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063989.

Fetzer:1981:PCC

[146] James H. Fetzer and Donald E. Nute. A probabilistic causal calculus: Conflicting conceptions. Synthese, 48(3):493, September 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063990;   http://link.
springer.com/content/pdf/10.1007/BF01063990.pdf.

Adams:1981:AT

[147] Robert Merrihew Adams. Actualism and thisness. Synthese, 49(1):3– 41, October 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063914.

Schiffer:1981:ITR

[148] Stephen Schiffer. Indexicals and the theory of reference. Synthese, 49 (1):43–100, October 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01063915.

Smith:1981:ISR

[149] David Woodruff Smith. Indexical sense and reference. Synthese, 49(1): 101–127, October 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01063916.

Stalnaker:1981:IB

[150] Robert C. Stalnaker. Indexical belief. Synthese, 49(1):129–151, October 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01063917.

McGinn:1981:MR

[151] Colin McGinn. The mechanism of reference. Synthese, 49(2):157–186, November 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064297.

Peacocke:1981:DTP

[152] Christopher Peacocke. Demonstrative thought and psychological explanation. Synthese, 49(2):187–217, November 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064298.

Bach:1981:RA

[153] Kent Bach. Referential/ attributive. Synthese, 49(2):219–244, November 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF01064299.

Baker:1981:MAD

[154] Lynne Rudder Baker. On making and attributing demonstrative reference. Synthese, 49(2):245–273, November 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064300.

Castaneda:1981:SPI

[155] Hector-Neri Castañeda. The semiotic profile of indexical (experiential) reference. Synthese, 49(2):275–316, November 1981. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064301.

Maloney:1981:NWE

[156] J. Christopher Maloney. A new way up from empirical foundations. Synthese, 49(3):317–335, December 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064125.

Bradie:1981:ACE

[157] Michael Bradie. Adequacy conditions and event identity. Synthese, 49(3): 337–374, December 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064126.

Hung:1981:TCL

[158] H.-C. Hung. Theories, catalogues, and languages. Synthese, 49(3):375– 394, December 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064127.

Horgan:1981:TPS

[159] Terence Horgan. Token physicalism, supervenience, and the generality of physics. Synthese, 49(3):395–413, December 1981. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064128.

Packard:1981:POS

[160] Dennis J. Packard. Plausibility orderings and social choice. Synthese, 49 (3):415–418, December 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064129.

Wettersten:1981:TNJ

[161] John Wettersten. On two non-justificationist moves. Synthese, 49(3): 419–421, December 1981. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064130.

Stenius:1981:SI

[162] Erik Stenius. Sets II. Synthese, 49(3):423–426, December 1981. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064131.

Slezak:1981:LPR

[163] Peter Slezak. Language and psychological reality: Discussion of Rudolf Botha’s study. Synthese, 49(3):427–440, December 1981. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064132. See [23] and reply [230].

Plendl:1982:EP

[164] Hans S. Plendl. Editorial preface. Synthese, 50(1):3, January 1982. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413722.

Cushing:1982:MMC

[165] James T. Cushing. Models and methodologies in current theoretical highenergy physics. Synthese, 50(1):5–101, January 1982. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413723.

Shrader-Frechette:1982:CCE

[166] K. S. Shrader-Frechette. Comments on Cushing’s essay. Synthese, 50 (1):103–108, January 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413724.

Cushing:1982:R

[167] James T. Cushing. A response. Synthese, 50(1):109–123, January 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413725.

Shrader-Frechette:1982:QQN

[168] K. S. Shrader-Frechette. Quark quantum numbers and the problem of microphysical observation. Synthese, 50(1):125–145, January 1982. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413726.

Albright:1982:CCV

[169] J. R. Albright. Comments concerning the visual acuity of quark hunters. Synthese, 50(1):147–152, January 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413727.

Shrader-Frechette:1982:CVA

[170] K. S. Shrader-Frechette. Consensus and the visual acuity of quark huntersa response. Synthese, 50(1):153–155, January 1982. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413728.

Gruender:1982:OQ

[171] David Gruender. On observing quarks. Synthese, 50(1):157–162, January 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413729.

Plendl:1982:EPPa

[172] Hans S. Plendl. Editorial preface to part II. Synthese, 50(2):165, February 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00416900.

Bruer:1982:CLQ

[173] John T. Bruer. The classical limit of quantum theory. Synthese, 50(2): 167–212, February 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00416901.

Weingard:1982:SS

[174] Robert Weingard and Gerrit Smith. Spin and space. Synthese, 50(2): 213–231, February 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00416902.

Ehrlich:1982:NIH

[175] Philip Ehrlich. Negative, infinite, and hotter than infinite temperatures. Synthese, 50(2):233–277, February 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00416903.

Fine:1982:SLM

[176] Arthur Fine. Some local models for correlation experiments. Synthese, 50 (2):279–294, February 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00416904.

Anonymous:1982:Aa

[177] Anonymous. Announcement. Synthese, 50(2):295, February 1982. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00416905.

Plendl:1982:EPPb

[178] Hans S. Plendl. Editorial preface to part III. Synthese, 50(3):299, March 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413886.

Urani:1982:ESR

[179] John Urani and George Gale. An extension of special relativity to accelerating frames and some of its philosophical implications. Synthese, 50 (3):301–323, March 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413887.

Fraisse:1982:QAF

[180] Roland Fraïssé. Quelques arguments en faveur de l’interprétation de la mécanique quantique par la ramification d’Everett. (French) [Some arguments in favor of the interpretation of quantum mechanics by the Everett model]. Synthese, 50(3):325–357, March 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413888.

Melvin:1982:TUF

[181] Mael A. Melvin. Towards unified field theory: Quantitative differences and qualitative sameness. Synthese, 50(3):359–397, March 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413889.

Finkelstein:1982:CC

[182] David Finkelstein. Cosmological choices. Synthese, 50(3):399–420, March 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413890.

Levy:1982:CRE

[183] Edwin Levy. Causal-relevance explanation: Salmon’s theory and its relation to Reichenbach. Synthese, 50(3):423–445, March 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413891.

Anonymous:1982:IDB

[184] Anonymous. IUHPS/DLMPS Bulletin No. 6. Synthese, 50(3):446– 447, March 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413892.

Macmillan:1982:EP

[185] C. J. B. Macmillan. Editorial preface. Synthese, 51(1):3, April 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413845.

Martin:1982:TDC

[186] Jane Roland Martin. Two dogmas of curriculum. Synthese, 51(1): 5–20, April 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413846.

Harrah:1982:WSW

[187] David Harrah. What should we teach about questions? Synthese, 51 (1):21–38, April 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413847.

Hintikka:1982:DMT

[188] Jaakko Hintikka. A dialogical model of teaching. Synthese, 51(1):39– 59, April 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413848.

Ennis:1982:IIA

[189] Robert H. Ennis. Identifying implicit assumptions. Synthese, 51(1): 61–86, April 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413849.

McClellan:1982:CLO

[190] James E. McClellan. The concept of learning: Once more with (logical) expression. Synthese, 51(1):87–116, April 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413850.

Airaksinen:1982:MED

[191] Timo Airaksinen. Moral education and democracy in the school. Synthese, 51(1):117–134, April 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413851.

Anonymous:1982:Ab

[192] Anonymous. Announcement. Synthese, 51(1):135–136, April 1982. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413852.

Anonymous:1982:Ea

[193] Anonymous. Errata. Synthese, 51(1):137, April 1982. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413853;   http://link.
springer.com/content/pdf/10.1007/BF00413853.pdf.

Putnam:1982:WTI

[194] Hilary Putnam. Why there isn’t a ready-made world. Synthese, 51(2): 141–167, May 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413825.

Glymour:1982:CSC

[195] Clark Glymour. Conceptual scheming or confessions of a metaphysical realist. Synthese, 51(2):169–180, May 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413826.

Horwich:1982:TFR

[196] Paul Horwich. Three forms of realism. Synthese, 51(2):181–201, May 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413827.

Wright:1982:SF

[197] Crispin Wright. Strict finitism. Synthese, 51(2):203–282, May 1982. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413828.

Clendinnen:1982:R

[198] F. John Clendinnen. Review. Synthese, 51(2):283–291, May 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413829.

Bell:1982:CTS

[199] J. L. Bell. Categories, toposes and sets. Synthese, 51(3):293–337, June 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485258.

Stemmer:1982:SLP

[200] Nathan Stemmer. A solution to the lottery paradox. Synthese, 51(3): 339–353, June 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485259.

Oddie:1982:CVN

[201] G. J. Oddie. Cohen on verisimilitude and natural necessity. Synthese, 51(3):355–379, June 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485260.

Wassermann:1982:LN

[202] Gerhard D. Wassermann. On the laws of nature. Synthese, 51(3):381– 396, June 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485261.

Macbeath:1982:WWD

[203] Murray Macbeath. Who was Dr Who’s father? Synthese, 51(3):397– 430, June 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485262.

vanBenthem:1982:LSS

[204] Johan van Benthem. The logical study of science. Synthese, 51(3): 431–472, June 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485263.

Putnam:1982:WRC

[205] Hilary Putnam. Why reason can’t be naturalized. Synthese, 52(1): 3–23, July 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485252.

VanFraassen:1982:CRE

[206] Bas C. Van Fraassen. The Charybdis of realism: Epistemological implications of Bell’s Inequality. Synthese, 52(1):25–38, July 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485253. See comment [437].

Pearce:1982:RFS

[207] David Pearce and Veikko Rantala. Realism and formal semantics. Synthese, 52(1):39–53, July 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485254.

Dummett:1982:R

[208] Michael Dummett. Realism. Synthese, 52(1):55–112, July 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485255.

McGinn:1982:RSC

[209] Colin McGinn. Realist semantics and content-ascription. Synthese, 52(1): 113–134, July 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485256.

Bell:1982:CN

[210] Martin Bell. Critical notice. Synthese, 52(1):135–166, July 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485257.

Hylton:1982:AIT

[211] Peter Hylton. Analyticity and the indeterminacy of translation. Synthese, 52(2):167–184, August 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869192.

Burdick:1982:LFP

[212] Howard Burdick. A logical form for the propositional attitudes. Synthese, 52(2):185–230, August 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869193.

Quine:1982:BA

[213] W. V. Quine. Burdick’s attitudes. Synthese, 52(2):231–232, August 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869194.

DAgostino:1982:SLA

[214] F. B. D’Agostino and H. R. Burdick. Symbolism and literalism in anthropology. Synthese, 52(2):233–265, August 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869195.

Bartlett:1982:RCC

[215] Steven J. Bartlett. Referential consistency as a a criterion of meaning. Synthese, 52(2):267–282, August 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869196.

Shanon:1982:CSV

[216] Benny Shanon. A calculus of semantic values. Synthese, 52(2):283– 298, August 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869197.

Karmo:1982:OPO

[217] Toomas Karmo. Occurrences and pseudo-occurrences. Synthese, 52(2): 299–312, August 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869198.

Roper:1982:ML

[218] James E. Roper. Models and lawlikeness. Synthese, 52(2):313–323, August 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869199.

Parsons:1982:RA

[219] Charles Parsons. Review article. Synthese, 52(2):325–343, August 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869200.

Sintonen:1982:RU

[220] Matti Sintonen. Realism and understanding. Synthese, 52(3):347–378, September 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869561.

Rasmussen:1982:RL

[221] Stig Alstrup Rasmussen and Jens Ravnkilde. Realism and logic. Synthese, 52(3):379–437, September 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869562.

Pearce:1982:RR

[222] David Pearce and Veikko Rantala. Realism and reference. Synthese, 52(3): 439–448, September 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869563.

Daum:1982:SEC

[223] Adam Daum. Schlick’s empiricist critical realism. Synthese, 52(3):449– 493, September 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869564.

McCabe:1982:DPU

[224] Viki McCabe. The direct perception of universals: A theory of knowledge acquisition. Synthese, 52(3):495–513, September 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:

//link.springer.com/article/10.1007/BF00869565.

Choy:1982:MBR

[225] Victoria Choy. Mind-body, realism and Rorty’s therapy. Synthese, 52(3): 515–541, September 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869566.

Anonymous:1982:Eb

[226] Anonymous. Errata. Synthese, 52(3):543–544, September 1982. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869567;   http:
//link.springer.com/content/pdf/10.1007/BF00869567.pdf.

Smart:1982:PBE

[227] J. J. C. Smart. Prior and the basis of ethics. Synthese, 53(1):3–17, October 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00500110.

Hughes:1982:SSO

[228] G. E. Hughes. Some strong omnitemporal logics. Synthese, 53(1):19– 42, October 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00500111.

Fine:1982:FOM

[229] Kit Fine. First-order modal theories III — facts. Synthese, 53(1):43– 122, October 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00500112.

Botha:1982:CMR

[230] Rudolf P. Botha. On Chomskyan mentalism: A reply to Peter Slezak. Synthese, 53(1):123–141, October 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00500113. See [23, 163].

Jelinek:1982:R

[231] Eloise Jelinek. Review. Synthese, 53(1):143–153, October 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00500114.

Anonymous:1982:Ac

[232] Anonymous. Announcement. Synthese, 53(1):I, October 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00500115.

Ross:1982:P

[233] Stephanie A. Ross and Paul A. Roth. Preface. Synthese, 53(2):157– 158, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484894.

Dennett:1982:HSH

[234] Daniel C. Dennett. How to study human consciousness empirically or nothing comes to mind. Synthese, 53(2):159–180, November 1982. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484895.

Rorty:1982:CD

[235] Richard Rorty. Comments on Dennett. Synthese, 53(2):181–187, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484896.

Hofstadter:1982:WSW

[236] Douglas R. Hofstadter. Who shoves whom around inside the careenium? or what is the meaning of the word “I”? Synthese, 53(2):189– 218, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484897.

Steele:1982:CH

[237] Guy L. Steele, Jr. Comments on Hofstadter. Synthese, 53(2):219–226, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484898.

Parfit:1982:PIR

[238] Derek Parfit. Personal identity and rationality. Synthese, 53(2):227– 241, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484899.

Regan:1982:CP

[239] Donald Regan. Comments on parfit. Synthese, 53(2):243–249, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484900.

Anonymous:1982:SD

[240] Anonymous. Summary of discussion. Synthese, 53(2):251–256, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484901.

Frankfurt:1982:IWW

[241] Harry Frankfurt. The importance of what we care about. Synthese, 53 (2):257–272, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484902.

Baier:1982:CAC

[242] Annette C. Baier. Caring about caring: A reply to Frankfurt. Synthese, 53 (2):273–290, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484903.

Macintyre:1982:CF

[243] Alasdair Macintyre. Comments on Frankfurt. Synthese, 53(2):291–294, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484904.

Macintyre:1982:HMA

[244] Alasdair Macintyre. How moral agents became ghosts or why the history of ethics diverged from that of the philosophy of mind. Synthese, 53(2): 295–312, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484905.

Dworkin:1982:RM

[245] Gerald Dworkin. Reply to MacIntyre. Synthese, 53(2):313–318, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484906.

Frankfurt:1982:CM

[246] Harry Frankfurt. Comments on MacIntyre. Synthese, 53(2):319–321, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484907.

Rorty:1982:CPM

[247] Richard Rorty. Contemporary philosophy of mind. Synthese, 53(2):323– 348, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF00484908. See comments [248].

Dennett:1982:CR

[248] Daniel C. Dennett. Comments on Rorty. Synthese, 53(2):349–356, November 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484909. See [247].

Gardenfors:1982:UPR

[249] Peter Gärdenfors and Nils-Eric Sahlin. Unreliable probabilities, risk taking, and decision making. Synthese, 53(3):361–386, December 1982. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486156.

Levi:1982:IPR

[250] Isaac Levi. Ignorance, probability and rational choice. Synthese, 53(3): 387–417, December 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486157.

VonPlato:1982:SED

[251] Jan Von Plato. The significance of the ergodic decomposition of stationary measures for the interpretation of probability. Synthese, 53(3): 419–432, December 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486158.

Gardenfors:1982:RL

[252] Peter Gärdenfors and Nils-Eric Sahlin. Reply to Levi. Synthese, 53(3): 433–438, December 1982. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486159.

Rehder:1982:CPC

[253] Wulf Rehder. Conditions for probabilities of conditionals to be conditional probabilities. Synthese, 53(3):439–443, December 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00486160.

Hellman:1982:EBS

[254] Geoffrey Hellman. Einstein and Bell: Strengthening the case for microphysical randomness. Synthese, 53(3):445–460, December 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00486161.

Hellman:1982:SEL

[255] Geoffrey Hellman. Stochastic Einstein-locality and the Bell theorems. Synthese, 53(3):461–503, December 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486162.

Hellman:1982:C

[256] Geoffrey Hellman. Corrigendum. Synthese, 53(3):504, December 1982. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486163;   http:
//link.springer.com/content/pdf/10.1007/BF00486163.pdf.

Anonymous:1982:Ad

[257] Anonymous. Announcement. Synthese, 53(3):505, December 1982. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486164.

Vaina:1983:SMO

[258] Lucia Vaina. From shapes and movements to objects and actions. Synthese, 54(1):3–36, January 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869461.

Mcgilvray:1983:C

[259] James A. Mcgilvray. To color. Synthese, 54(1):37–70, January 1983. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869462.

Topper:1983:ARO

[260] David R. Topper. Art in the realist ontology of J. J. Gibson. Synthese, 54(1):71–83, January 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869463.

Reed:1983:TTI

[261] Edward S. Reed. Two theories of the intentionality of perceiving. Synthese, 54(1):85–94, January 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869464.

Smith:1983:DSB

[262] David Woodruff Smith. Is this a dagger I see before me? Synthese, 54 (1):95–114, January 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869465.

Saarinen:1983:LPS

[263] Esa Saarinen. On the logic of perception sentences. Synthese, 54(1):115– 128, January 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869466.

Vlach:1983:SSP

[264] Frank Vlach. On situation semantics for perception. Synthese, 54(1):129– 152, January 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869467.

Hintikka:1983:SPW

[265] Jaakko Hintikka. Situations, possible worlds, and attitudes. Synthese, 54 (1):153–162, January 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869468.

Hintikka:1983:Pa

[266] Jaakko Hintikka. Preface. Synthese, 54(2):165, February 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869389.

Lepore:1983:WMT

[267] Ernest Lepore. What model theoretic semantics cannot do? Synthese, 54 (2):167–187, February 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869390.

Bach:1983:RWR

[268] Kent Bach. Russell was right (almost). Synthese, 54(2):189–207, February 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869391.

Richards:1983:ADD

[269] Barry Richards. Anaphora, descriptions and discourse representations. Synthese, 54(2):209–233, February 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869392.

Forbes:1983:TV

[270] Graeme Forbes. Thisness and vagueness. Synthese, 54(2):235–259, February 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869393.

Smith:1983:K

[271] David Woodruff Smith. Kantifying in. Synthese, 54(2):261–273, February 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869394.

Kraut:1983:TND

[272] Robert Kraut. There are no de dicto attitudes. Synthese, 54(2):275– 294, February 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869395.

Loewer:1983:DDD

[273] Barry Loewer and Marvin Belzer. Dyadic deontic detachment. Synthese, 54(2):295–318, February 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869396.

Anonymous:1983:A

[274] Anonymous. Announcement. Synthese, 54(2):319, February 1983. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869397.

Hintikka:1983:Pb

[275] Jaakko Hintikka. Preface. Synthese, 54(3):323, March 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00877684.

Harnish:1983:PD

[276] Robert M. Harnish. Pragmatic derivations. Synthese, 54(3):325– 373, March 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00877685.

Gilbert:1983:ACL

[277] Margaret Gilbert. Agreements, conventions, and language. Synthese, 54 (3):375–407, March 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00877686.

Dancy:1983:AE

[278] R. M. Dancy. Aristotle and existence. Synthese, 54(3):409–442, March 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00877687.

Hintikka:1983:SGA

[279] Jaakko Hintikka. Semantical games, the alleged ambiguity of ‘is’, and Aristotelian categories. Synthese, 54(3):443–468, March 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00877688.

Bach:1983:R

[280] Kent Bach and Robert M. Harnish. Review. Synthese, 54(3):469– 493, March 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00877689.

Anonymous:1983:IDB

[281] Anonymous. IUHPS/DLMPS Bulletin No. 7. Synthese, 54(3):495– 496, March 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00877690.

Audi:1983:EPa

[282] Robert Audi. Editorial preface. Synthese, 55(1):1, April 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485370.

Dretske:1983:EB

[283] Fred I. Dretske. The epistemology of belief. Synthese, 55(1):3–19, April 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485371.

Goldman:1983:ETP

[284] Alvin I. Goldman. Epistemology and the theory of problem solving. Synthese, 55(1):21–48, April 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485372.

Sosa:1983:NUE

[285] Ernest Sosa. Nature unmirrored, epistemology naturalized. Synthese, 55 (1):49–72, April 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485373.

Alston:1983:WWI

[286] William P. Alston. What’s wrong with immediate knowledge? Synthese, 55(1):73–95, April 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485374.

Shatz:1983:FCL

[287] David Shatz. Foundationalism, coherentism, and the levels gambit. Synthese, 55(1):97–118, April 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485375.

Audi:1983:FED

[288] Robert Audi. Foundationalism, epistemic dependence, and defeasibility. Synthese, 55(1):119–139, April 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485376.

Audi:1983:EPb

[289] Robert Audi. Editorial preface. Synthese, 55(2):141, May 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485064.

Klein:1983:RK

[290] Peter D. Klein. Real knowledge. Synthese, 55(2):143–164, May 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485065.

Hilpinen:1983:SJ

[291] Risto Hilpinen. Skepticism and justification. Synthese, 55(2):165– 173, May 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485066.

Yourgrau:1983:KRA

[292] Palle Yourgrau. Knowledge and relevant alternatives. Synthese, 55(2): 175–190, May 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485067.

Lehrer:1983:JTC

[293] Keith Lehrer and Stewart Cohen. Justification, truth, and coherence. Synthese, 55(2):191–207, May 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485068.

Schmitt:1983:KJR

[294] Frederick F. Schmitt. Knowledge, justification, and reliability. Synthese, 55(2):209–229, May 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485069.

Pollock:1983:EP

[295] John L. Pollock. Epistemology and probability. Synthese, 55(2):231– 252, May 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485070.

Pappas:1983:OK

[296] George S. Pappas. Ongoing knowledge. Synthese, 55(2):253–267, May 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485071.

Naylor:1983:JMK

[297] Andrew Naylor. Justification in memory knowledge. Synthese, 55(2): 269–286, May 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485072.

Falmagne:1983:SMQ

[298] J. C. Falmagne and L. Narens. Scales and meaningfulness of quantitative laws. Synthese, 55(3):287–325, June 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485824.

Lenker:1983:NOT

[299] Terry D. Lenker and Richard St. André. Near orderings of topological spaces. Synthese, 55(3):327–331, June 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485825.

Heiner:1983:MAP

[300] Ronald A. Heiner and Dennis J. Packard. More about plausibility orderings. Synthese, 55(3):333–337, June 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485826.

Lehrer:1983:PAI

[301] Keith Lehrer and Carl Wagner. Probability amalgamation and the independence issue: A reply to Laddaga. Synthese, 55(3):339–346, June 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485827. See [1].

Sowden:1983:TDP

[302] Lanning Sowden. That there is a dilemma in the prisoners’ dilemma. Synthese, 55(3):347–352, June 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485828.

Mortensen:1983:RV

[303] Chris Mortensen. Relevance and verisimilitude. Synthese, 55(3):353– 364, June 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485829.

Burdick:1983:SLA

[304] Howard Burdick. On symbolism and literalism in anthropology. Synthese, 55(3):365–371, June 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485830.

Bogen:1983:R

[305] James Bogen. Review. Synthese, 55(3):373–388, June 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485831.

Pearce:1983:NFM

[306] David Pearce and Veikko Rantala. New foundations for metascience. Synthese, 56(1):1–26, July 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485393.

Macbeath:1983:CTR

[307] Murray Macbeath. Communication and time reversal. Synthese, 56(1): 27–46, July 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485394.

Stairs:1983:LPQ

[308] Allen Stairs. On the logic of pairs of quantum systems. Synthese, 56 (1):47–60, July 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485395.

Siegel:1983:BEN

[309] Harvey Siegel. Brown on epistemology and the new philosophy of science. Synthese, 56(1):61–89, July 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485396.

Brown:1983:RS

[310] Harold I. Brown. Response to Siegel. Synthese, 56(1):91–105, July 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485397.

Rynasiewicz:1983:R

[311] Robert Rynasiewicz. Review. Synthese, 56(1):107–120, July 1983. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485398.

Hintikka:1983:Pc

[312] Jaakko Hintikka. Preface. Synthese, 56(2):121, August 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485319.

Sluga:1983:ST

[313] Hans Sluga. Subjectivity in the Tractatus. Synthese, 56(2):123–139, August 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485320.

Fogelin:1983:WI

[314] Robert J. Fogelin. Wittgenstein on identity. Synthese, 56(2):141– 154, August 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485321.

Hintikka:1983:SRW

[315] Jaakko Hintikka and Merrill B. Hintikka. Some remarks on (Wittgensteinian) logical form. Synthese, 56(2):155–170, August 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485322.

Maury:1983:RLF

[316] André Maury. Reality and logical form. Synthese, 56(2):171–180, August 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485323.

Arrington:1983:RWT

[317] Robert L. Arrington. Representation in Wittgenstein’s Tractatus and middle writings. Synthese, 56(2):181–198, August 1983. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485324.

Wiesenthal:1983:VSP

[318] L. Wiesenthal. Visual space from the perspective of possible-worlds semantics, I. Synthese, 56(2):199–238, August 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485325.

Gottlieb:1983:WCT

[319] Diane F. Gottlieb. Wittgenstein’s critique of the Tractatus view of rules. Synthese, 56(2):239–251, August 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485326.

Wavell:1983:WDU

[320] Bruce B. Wavell. Wittgenstein’s doctrine of use. Synthese, 56(3):253– 264, September 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485465.

Goldfarb:1983:WYB

[321] Warren D. Goldfarb. I want you to bring me a slab: Remarks on the opening sections of the philosophical investigations. Synthese, 56(3):265– 282, September 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485466.

Dancy:1983:AC

[322] R. M. Dancy. Alien concepts. Synthese, 56(3):283–300, September 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485467.

Gilbert:1983:QWL

[323] Margaret Gilbert. On the question whether language has a social nature: Some aspects of winch and others on Wittgenstein. Synthese, 56(3):301– 318, September 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485468.

Luckhardt:1983:WB

[324] C. Grant Luckhardt. Wittgenstein and behaviorism. Synthese, 56(3): 319–338, September 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485469.

Obermeier:1983:WLA

[325] Klaus K. Obermeier. Wittgenstein on language and artificial intelligence: The Chinese-room thought experiment revisited. Synthese, 56(3):339– 349, September 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485470.

Ziff:1983:RMR

[326] Paul Ziff. Remarks on Mittgenstein’s remarks on the foundations of mathematics. Synthese, 56(3):351–361, September 1983. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485471.

Macmillan:1983:CI

[327] C. J. B. Macmillan. On certainty and indoctrination. Synthese, 56(3): 363–372, September 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485472.

Black:1983:OK

[328] Carolyn Black. Obvious knowledge. Synthese, 56(3):373–385, September 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485473.

Bowen:1983:CCP

[329] Paul D. Bowen. Causation in classical physics. Synthese, 57(1):1– 20, October 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064064.

Hauptli:1983:QTA

[330] Bruce Hauptli. Quine’s theorizing about theories. Synthese, 57(1):21– 33, October 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064065.

Falmagne:1983:RUM

[331] Jean-Claude Falmagne. A random utility model for a belief function. Synthese, 57(1):35–48, October 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064066.

Priest:1983:ARA

[332] Graham Priest. An anti-realist account of mathematical truth. Synthese, 57(1):49–65, October 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064067.

Ellett:1983:LCM

[333] Frederick S. Ellett, Jr. and David P. Ericson. The logic of causal methods in social science. Synthese, 57(1):67–82, October 1983. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064068.

Mcdonell:1983:PUS

[334] Neil Mcdonell. Are pictures unavoidably specific? Synthese, 57(1):83– 98, October 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064069.

Friedman:1983:SPO

[335] Joel I. Friedman. Spinoza’s problem of “other minds”. Synthese, 57(1): 99–126, October 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064070.

F:1983:PP

[336] J. H. F. Preface to Part I. Synthese, 57(2):127, November 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01063997.

Smokler:1983:IRC

[337] Howard Smokler. Institutional rationality: The complex norms of science. Synthese, 57(2):129–138, November 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01063998.

Tweney:1983:RPI

[338] Ryan D. Tweney and Michael E. Doherty. Rationality and the psychology of inference. Synthese, 57(2):139–161, November 1983. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01063999.

Cherniak:1983:RSH

[339] Christopher Cherniak. Rationality and the structure of human memory. Synthese, 57(2):163–186, November 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01064000.

Mullane:1983:DDR

[340] Harvey Mullane. Defense, dreams and rationality. Synthese, 57(2):187– 204, November 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064001.

Macklin:1983:PCR

[341] Ruth Macklin. Philosophical conceptions of rationality and psychiatric notions of competency. Synthese, 57(2):205–224, November 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064002.

Gewirth:1983:RR

[342] Alan Gewirth. The rationality of reasonableness. Synthese, 57(2):225– 247, November 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064003.

Thomas:1983:RMA

[343] Laurence Thomas. Rationality and moral autonomy: An essay in moral psychology. Synthese, 57(2):249–266, November 1983. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064004.

F:1983:PPI

[344] J. H. F. Preface to part II. Synthese, 57(3):267, December 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064697.

Backman:1983:PSR

[345] Wayne Backman. Practical and scientific rationality: A difficulty for Levi’s epistemology. Synthese, 57(3):269–276, December 1983. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF01064698.

Kaplan:1983:PSR

[346] Mark Kaplan. Practical and scientific rationality: A Bayesian perspective on Levi’s difficulty. Synthese, 57(3):277–282, December 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064699.

Lehrer:1983:RWA

[347] Keith Lehrer. Rationality as weighted averaging. Synthese, 57(3):283– 295, December 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064700.

Mayo:1983:OTS

[348] Deborah G. Mayo. An objective theory of statistical testing. Synthese, 57 (3):297–340, December 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.

1007/BF01064701. See response [597].

Harsanyi:1983:BDT

[349] John C. Harsanyi. Bayesian decision theory, subjective and objective probabilities, and acceptance of empirical hypotheses. Synthese, 57(3): 341–365, December 1983. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF01064702.

Fetzer:1983:POD

[350] James H. Fetzer. Probability and objectivity in deterministic and indeterministic situations. Synthese, 57(3):367–386, December 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01064703.

Eells:1983:OPT

[351] Ellery Eells. Objective probability theory theory. Synthese, 57(3):387– 442, December 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01064704.

Fenstad:1983:IUH

[352] Jens Erik Fenstad. International Union of the History and Philosophy of Science Division of Logic, Methodology and Philosophy of Science Bulletin No. 8. Synthese, 57(3):443–453, December 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01064705.

Almog:1984:WYB

[353] Joseph Almog. Would you believe that? Synthese, 58(1):1–37, January 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485359.

Suppes:1984:PAR

[354] Patrick Suppes. A puzzle about responses and congruence of meaning. Synthese, 58(1):39–50, January 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485360. See rejoinder [355].

Almog:1984:BII

[355] Joseph Almog. Believe it or not: It is a puzzle. Rejoinder to Suppes. Synthese, 58(1):51–61, January 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485361. See [354].

Wettstein:1984:HBG

[356] Howard K. Wettstein. How to bridge the gap between meaning and reference. Synthese, 58(1):63–84, January 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485362.

DAgostino:1984:CC

[357] Fred D’Agostino. Chomsky on creativity. Synthese, 58(1):85–117, January 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485363.

Dancy:1984:EI

[358] Jonathan Dancy. Even-ifs. Synthese, 58(2):119–128, February 1984. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF03055302.

Williams:1984:LLR

[359] Meredith Williams. Language learning and the representational theory of mind. Synthese, 58(2):129–151, February 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF03055303.

Kittay:1984:IM

[360] Eva Feder Kittay. The identification of metaphor. Synthese, 58(2):153– 202, February 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF03055304.

Bertolet:1984:INF

[361] Rod Bertolet. Inferences, names, and fictions. Synthese, 58(2):203– 218, February 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF03055305.

Rolf:1984:S

[362] Bertil Rolf. Sorites. Synthese, 58(2):219–250, February 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF03055306.

Cartwright:1984:PPN

[363] Helen Morris Cartwright. Parts and partitives: Notes on what things are made of. Synthese, 58(2):251–277, February 1984. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF03055307.

Anonymous:1984:Aa

[364] Anonymous. Announcement. Synthese, 58(2):278–279, February 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF03055308.

Blackburn:1984:ISB

[365] Simon Blackburn. The individual strikes back. Synthese, 58(3):281– 301, March 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485244.

Budd:1984:WMI

[366] Malcolm Budd. Wittgenstein on meaning, interpretation and rules. Synthese, 58(3):303–323, March 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485245.

McDowell:1984:WFR

[367] John McDowell. Wittgenstein on following a rule. Synthese, 58(3):325– 363, March 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485246.

Peacocke:1984:CCC

[368] Christopher Peacocke. Colour concepts and colour experience. Synthese, 58(3):365–381, March 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485247.

Wright:1984:STA

[369] Crispin Wright. Second thoughts about criteria. Synthese, 58(3):383– 405, March 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485248.

Baker:1984:MWK

[370] G. P. Baker and P. M. S. Hacker. On misunderstanding Wittgenstein: Kripke’s private language argument. Synthese, 58(3):407–450, March 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485249.

Carruthers:1984:BHW

[371] Peter Carruthers. Baker and Hacker’s Wittgenstein. Synthese, 58(3): 451–479, March 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485250.

Anonymous:1984:Ab

[372] Anonymous. Announcement. Synthese, 58(3):480, March 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485251.

Floistad:1984:P

[373] Guttorm Fløistad and Dagfinn Føllesdal. Preface. Synthese, 59(1):1, April 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00873277.

Davidson:1984:CC

[374] Donald Davidson. Communication and convention. Synthese, 59(1): 3–17, April 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873278.

Apel:1984:CD

[375] Karl-Otto Apel. Comments on Davidson. Synthese, 59(1):19–26, April 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00873279.

Hintikka:1984:HYL

[376] Jaakko Hintikka. A hundred years later: The rise and fall of Frege’s influence in language theory. Synthese, 59(1):27–49, April 1984. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00873280. See com-

ments [377].

Guenthner:1984:CHH

[377] Franz Guenthner. Comments on Hintikka’s ‘A hundred years later’. Synthese, 59(1):51–58, April 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00873281. See [376].

Ladriere:1984:SSF

[378] Jean Ladrière. Signification et signifiance. (French) [Meaning and meaningfulness]. Synthese, 59(1):59–67, April 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00873282.

Markovic:1984:LI

[379] Mihailo Markovic. The language of ideology. Synthese, 59(1):69–88, April 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00873283.

Black:1984:RAP

[380] Max Black. The radical ambiguity of a poem. Synthese, 59(1):89– 107, April 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873284.

Granger:1984:LIF

[381] Gilles G. Granger. Langage et individuation. (French) [Language and individuation]. Synthese, 59(1):109–114, April 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00873285.

Anonymous:1984:Ac

[382] Anonymous. Announcement. Synthese, 59(1):115, April 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00873286.

Hornstein:1984:IQN

[383] Norbert Hornstein. Interpreting quantification in natural language. Synthese, 59(2):117–150, May 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869415.

Eberle:1984:LRT

[384] Rolf A. Eberle. Logic with a relative truth predicate and “that”-terms. Synthese, 59(2):151–185, May 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869416.

Bonevac:1984:SCC

[385] Daniel Bonevac. Semantics for clausally complemented verbs. Synthese, 59(2):187–218, May 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869417.

Carleton:1984:PLU

[386] Lawrence R. Carleton. Programs, language understanding, and searle. Synthese, 59(2):219–230, May 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869418.

Fodor:1984:SWS

[387] J. A. Fodor. Semantics, Wisconsin style. Synthese, 59(3):231–250, June 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869335.

Maloney:1984:MML

[388] J. Christopher Maloney. The mundane mental language: How to do words with things. Synthese, 59(3):251–294, June 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869336.

Hill:1984:DTM

[389] Christopher S. Hill. In defense of type materialism. Synthese, 59(3): 295–320, June 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869337.

Horgan:1984:FTP

[390] Terence Horgan. Functionalism and token physicalism. Synthese, 59(3): 321–338, June 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869338.

Brennan:1984:S

[391] Andrew Brennan. Survival. Synthese, 59(3):339–361, June 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869339.

Linsky:1984:PQI

[392] Bernard Linsky. Phenomenal qualities and the identity of indistinguishables. Synthese, 59(3):363–380, June 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869340.

Tamny:1984:CCS

[393] Martin Tamny and Raphael Stern. City college studies in the history and philosophy of science and technology: Series editors’ preface. Synthese, 60(1):1–2, July 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485613.

Leblanc:1984:P

[394] Hugues Leblanc, Elliott Mendelson, and Alex Orenstein. Preface. Synthese, 60(1):3, July 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485614.

Fitting:1984:SAA

[395] Melvin Fitting. A symmetric approach to axiomatizing quantifiers and modalities. Synthese, 60(1):5–19, July 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485615.

Goodman:1984:KM

[396] Nicolas D. Goodman. The knowing mathematician. Synthese, 60(1): 21–38, July 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485616.

Gumb:1984:CKC

[397] Raymond D. Gumb. “Conservative” Kripke closures. Synthese, 60(1): 39–49, July 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485617.

Hiz:1984:FLI

[398] Henry Hiz. Frege, Leśniewski and information semantics on the resolution of antinomies. Synthese, 60(1):51–72, July 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485618.

Jeffrey:1984:FP

[399] Richard Jeffrey. De Finetti’s probabilism. Synthese, 60(1):73–90, July 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485619.

Leblanc:1984:PFT

[400] Hugues Leblanc and Charles G. Morgan. Probability functions and their assumption sets- the binary case. Synthese, 60(1):91–106, July 1984. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485620.

Harman:1984:LR

[401] Gilbert Harman. Logic and reasoning. Synthese, 60(1):107–127, July 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485621.

Myhill:1984:P

[402] John Myhill. Paradoxes. Synthese, 60(1):129–143, July 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485622.

Anonymous:1984:ENa

[403] Anonymous. Editorial note. Synthese, 60(1):144, July 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485623.

Orenstein:1984:RNS

[404] Alex Orenstein. Referential and nonreferential substitutional quantifiers. Synthese, 60(2):145–157, August 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485459.

Sieg:1984:FAP

[405] Wilfried Sieg. Foundations for analysis and proof theory. Synthese, 60 (2):159–200, August 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485460.

Smullyan:1984:CL

[406] Raymond M. Smullyan. Chameleonic languages. Synthese, 60(2):201– 224, August 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485461.

Stern:1984:RMS

[407] Raphael Stern. Relational model systems: The craft of logic. Synthese, 60(2):225–252, August 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485462.

Diller:1984:RIL

[408] J. Diller and A. S. Troelstra. Realizability and intuitionistic logic. Synthese, 60(2):253–282, August 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485463.

Anonymous:1984:Ad

[409] Anonymous. Announcement. Synthese, 60(2):283–284, August 1984. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485464.

Iglesias:1984:RTK

[410] Teresa Iglesias. Russell’s theory of knowledge and Wittgenstein’s earliest writings. Synthese, 60(3):285–332, September 1984. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485561.

Irvine:1984:RCS

[411] William B. Irvine. Russell’s construction of space from perspectives. Synthese, 60(3):333–347, September 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485562.

Dunlop:1984:WSS

[412] Charles E. M. Dunlop. Wittgenstein on sensation and ‘seeing-as’. Synthese, 60(3):349–367, September 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485563.

Mccarthy:1984:RIQ

[413] Timothy Mccarthy. Representation, intentionality, and quantifiers. Synthese, 60(3):369–411, September 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485564.

Bertolet:1984:RFF

[414] Rod Bertolet. Reference, fiction, and fictions. Synthese, 60(3):413–437, September 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485565.

Wettstein:1984:DGR

[415] Howard Wettstein. Did the Greeks really worship Zeus? Synthese, 60(3): 439–449, September 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485566.

Hintikka:1984:TNO

[416] Jaakko Hintikka. Are there nonexistent objects? Why not? But where are they? Synthese, 60(3):451–458, September 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485567.

Causey:1984:R

[417] Robert L. Causey. Review. Synthese, 60(3):459–466, September 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485568.

Anonymous:1984:ENb

[418] Anonymous. Editorial note. Synthese, 61(1):1, October 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485485.

Searle:1984:IPN

[419] John R. Searle. Intentionality and its place in nature. Synthese, 61(1): 3–16, October 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485486.

Mohanty:1984:ICH

[420] J. N. Mohanty. Intentionality, causality and holism. Synthese, 61(1): 17–33, October 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485487.

Miller:1984:PR

[421] Izchak Miller. Perceptual reference. Synthese, 61(1):35–59, October 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485488.

Smith:1984:CCP

[422] David Woodruff Smith. Content and context of perception. Synthese, 61(1):61–87, October 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485489.

Chisholm:1984:PI

[423] Roderick M. Chisholm. The primacy of the intentional. Synthese, 61 (1):89–109, October 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485490.

Tomberlin:1984:III

[424] James E. Tomberlin. Identity, intensionality, and intentionality. Synthese, 61(1):111–131, October 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485491.

Anonymous:1984:Ae

[425] Anonymous. Announcement. Synthese, 61(1):132, October 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485492.

Anonymous:1984:ENc

[426] Anonymous. Editorial note. Synthese, 61(2):133, November 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485311.

Castaneda:1984:TFG

[427] Hector-Neri Castañeda. Tomberlin, Frege, and Guise theory: A note on the methodology of dia-philosophical comparisons. Synthese, 61(2):135– 147, November 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485312.

Richards:1984:SCA

[428] William M. Richards. Self-consciousness and agency. Synthese, 61(2): 149–171, November 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485313.

Nelson:1984:NI

[429] R. J. Nelson. Naturalizing intentions. Synthese, 61(2):173–203, November 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485314.

Roth:1984:MNB

[430] Paul A. Roth. On missing Neurath’s boat: Some reflections on recent Quine literature. Synthese, 61(2):205–231, November 1984. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485315.

Thagard:1984:FKI

[431] Paul Thagard. Frames, knowledge, and inference. Synthese, 61(2):233– 259, November 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485316.

Adler:1984:R

[432] Jonathan E. Adler. Review. Synthese, 61(2):261–272, November 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485317.

Anonymous:1984:Af

[433] Anonymous. Announcements. Synthese, 61(2):273–274, November 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485318.

Hoy:1984:IIP

[434] Ronald C. Hoy. Inquiry, intrinsic properties, and the identity of indiscernibles. Synthese, 61(3):275–297, December 1984. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485055.

Price:1984:PPA

[435] Huw Price. The philosophy and physics of affecting the past. Synthese, 61 (3):299–323, December 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485056.

Ginsberg:1984:PQM

[436] Allen Ginsberg. On a paradox in quantum mechanics. Synthese, 61(3): 325–349, December 1984. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485057.

Stairs:1984:SCV

[437] Allen Stairs. Sailing into the Charybdis: Van Fraassen on Bell’s Theorem. Synthese, 61(3):351–359, December 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485058. See [206].

Kuipers:1984:ACO

[438] Theo A. F. Kuipers. An approximation of Carnap’s optimum estimation method. Synthese, 61(3):361–362, December 1984. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485059.

Fair:1984:PMT

[439] David Fair. Provability and mathematical truth. Synthese, 61(3):363– 385, December 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485060.

Clark:1984:H

[440] Peter Clark and Stephen Read. Hypertasks. Synthese, 61(3):387–390, December 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485061.

Thorpe:1984:SP

[441] Dale A. Thorpe. The sorites paradox. Synthese, 61(3):391–421, December 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485062.

Kroes:1984:OVM

[442] Peter Kroes. Objective versus minddependent theories of time flow. Synthese, 61(3):423–446, December 1984. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485063.

Loewer:1985:P

[443] Barry Loewer. Preface. Synthese, 62(1):1, January 1985. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485381.

Levi:1985:CSA

[444] Isaac Levi. Consensus as shared agreement and outcome of inquiry. Synthese, 62(1):3–11, January 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485382.

Nurmi:1985:SPL

[445] Hannu Nurmi. Some properties of the Lehrer–Wagner method for reaching rational consensus. Synthese, 62(1):13–24, January 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485383.

Schmitt:1985:CRW

[446] Frederick F. Schmitt. Consensus, respect, and weighted averaging. Synthese, 62(1):25–46, January 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485384.

Baird:1985:LWC

[447] Davis Baird. Lehrer/Wagner consensual probabilities do not adequately summarize the available information. Synthese, 62(1):47–62, January 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485385.

Kvanvig:1985:TUJ

[448] Jonathan L. Kvanvig. Is there an ‘us’ in ‘justification’? Synthese, 62 (1):63–73, January 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485386.

Forrest:1985:LWT

[449] Peter Forrest. The Lehrer/Wagner theory of consensus and the zero weight problem. Synthese, 62(1):75–78, January 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485387.

Loewer:1985:DC

[450] Barry Loewer and Robert Laddaga. Destroying the consensus. Synthese, 62(1):79–95, January 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485388.

Wagner:1985:FPW

[451] Carl Wagner. On the formal properties of weighted averaging as a method of aggregation. Synthese, 62(1):97–108, January 1985. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485389.

Lehrer:1985:CIO

[452] Keith Lehrer. Consensus and the ideal observer. Synthese, 62(1):109– 120, January 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485390.

Anonymous:1985:ANJ

[453] Anonymous. Announcing a new journal. Synthese, 62(1):121, January 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485391.

Anonymous:1985:Aa

[454] Anonymous. Announcement. Synthese, 62(1):122, January 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485392.

Anonymous:1985:EN

[455] Anonymous. Editorial note. Synthese, 62(2):123, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00486041.

Casari:1985:RFI

[456] Ettore Casari. Remarks on the foundational inquiry. Synthese, 62(2): 125–137, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486042.

Kreisel:1985:MLT

[457] Georg Kreisel. Mathematical logic: Tool and object lesson for science. Synthese, 62(2):139–151, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486043.

Prawitz:1985:RSA

[458] Dag Prawitz. Remarks on some approaches to the concept of logical consequence. Synthese, 62(2):153–171, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00486044.

Cellucci:1985:PTC

[459] Carlo Cellucci. Proof theory and complexity. Synthese, 62(2):173– 189, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486045.

Girard:1985:IL

[460] Jean-Yves Girard. Introduction to Π_{2}^{1}-logic. Synthese, 62(2):191–216, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486046.

Troelstra:1985:CSI

[461] A. S. Troelstra. Choice sequences and informal rigour. Synthese, 62(2): 217–227, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486047.

Feferman:1985:WF

[462] Solomon Feferman. Working foundations. Synthese, 62(2):229–254, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486048.

Takeuti:1985:PTS

[463] Gaisi Takeuti. Proof theory and set theory. Synthese, 62(2):255–263, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486049.

Magari:1985:SM

[464] Roberto Magari. The success of mathematics. Synthese, 62(2):265– 274, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486050.

Lolli:1985:FPC

[465] Gabriele Lolli. Foundational problems from computation theory. Synthese, 62(2):275–288, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486051.

Fenstad:1985:NAR

[466] Jens Erik Fenstad. Is nonstandard analysis relevant for the philosophy of mathematics? Synthese, 62(2):289–301, February 1985. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00486052.

Chiara:1985:SFP

[467] Maria Luisa Dalla Chiara. Some foundational problems in mathematics suggested by physics. Synthese, 62(2):303–315, February 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00486053.

Anonymous:1985:Ab

[468] Anonymous. Announcement. Synthese, 62(2):316, February 1985. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486054.

Creath:1985:TTS

[469] Richard Creath. Taking theories seriously. Synthese, 62(3):317–345, March 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869409.

Makinson:1985:HGI

[470] David Makinson. How to give it up: A survey of some formal aspects of the logic of theory change. Synthese, 62(3):347–363, March 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869410.

Kleiner:1985:IPS

[471] Scott A. Kleiner. Interrogatives, problems and scientific inquiry. Synthese, 62(3):365–428, March 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869411.

Finocchiaro:1985:ALH

[472] Maurice A. Finocchiaro. Aspects of the logic of history-of-science explanation. Synthese, 62(3):429–454, March 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869412.

Doppelt:1985:FRE

[473] Gerald Doppelt. Finocchiaro on rational explanation. Synthese, 62(3): 455–458, March 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869413.

Halbwachs:1985:GWM

[474] F. Halbwachs and A. Torunczyk. On Galileo’s writings on mechanics: An attempt at a semantic analysis of Viviani’s scholium. Synthese, 62 (3):459–484, March 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869414.

Skilling:1985:PP

[475] John Skilling. Prior probabilities. Synthese, 63(1):1–34, April 1985. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485953.

Shimony:1985:SPM

[476] Abner Shimony. The status of the principle of maximum entropy. Synthese, 63(1):35–53, April 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485954.

Skyrms:1985:MEI

[477] Brian Skyrms. Maximum entropy inference as a special case of conditionalization. Synthese, 63(1):55–74, April 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485955.

Domotor:1985:PKC

[478] Zoltan Domotor. Probability kinematics, conditionals, and entropy principles. Synthese, 63(1):75–114, April 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485956.

Jaynes:1985:SRO

[479] E. T. Jaynes. Some random observations. Synthese, 63(1):115–138, April 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485957.

Bar-Hillel:1985:GPP

[480] Maya Bar-Hillel and Avishai Margalit. Gideon’s paradox- a paradox of rationality. Synthese, 63(2):139–155, May 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485364.

Sorensen:1985:IVN

[481] Roy A. Sorensen. The iterated versions of Newcomb’s problem and the prisoner’s dilemma. Synthese, 63(2):157–166, May 1985. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485365.

Sobel:1985:CDC

[482] Jordan Howard Sobel. Circumstances and dominance in a causal decision theory. Synthese, 63(2):167–202, May 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485366.

Richter:1985:RGC

[483] Reed Richter. Rationality, group choice and expected utility. Synthese, 63(2):203–232, May 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485367.

Pitowsky:1985:SSI

[484] Itamar Pitowsky. On the status of statistical inferences. Synthese, 63(2): 233–247, May 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485368.

Aronson:1985:CVT

[485] Jerrold L. Aronson. Conditions versus transference: A reply to Ehring. Synthese, 63(2):249–257, May 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485369.

Walton:1985:NDL

[486] Douglas N. Walton. New directions in the logic of dialogue. Synthese, 63(3):259–274, June 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485596.

Harrah:1985:LMR

[487] David Harrah. A logic of message and reply. Synthese, 63(3):275–294, June 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485597.

Krabbe:1985:FSD

[488] Erik C. W. Krabbe. Formal systems of dialogue rules. Synthese, 63(3): 295–328, June 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485598.

Mackenzie:1985:NLB

[489] Jim Mackenzie. No logic before Friday. Synthese, 63(3):329–341, June 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485599.

Nowakowska:1985:FSD

[490] Maria Nowakowska. On a formal structure of a dialogue. Synthese, 63 (3):343–353, June 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485600.

Stump:1985:LDW

[491] Eleonore Stump. The logic of disputation in Walter Burley’s treatise on obligations. Synthese, 63(3):355–374, June 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485601.

Barth:1985:NFE

[492] E. M. Barth. A new field: Empirical logic bioprograms, logemes and logics as institutions. Synthese, 63(3):375–388, June 1985. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485602.

Machamer:1985:P

[493] Peter K. Machamer. Preface. Synthese, 64(1):1, July 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485709.

Sosa:1985:CVV

[494] Ernest Sosa. The coherence of virtue and the virtue of coherence. Synthese, 64(1):3–28, July 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485710.

Goldman:1985:RBE

[495] Alvin I. Goldman. The relation between epistemology and psychology. Synthese, 64(1):29–68, July 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485711.

Swain:1985:JRR

[496] Marshall Swain. Justification, reasons, and reliability. Synthese, 64(1): 69–92, July 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485712.

Nelson:1985:DP

[497] Jack Nelson. The diversity of perception. Synthese, 64(1):93–113, July 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485713.

Stich:1985:CMI

[498] Stephen P. Stich. Could man be an irrational animal? Synthese, 64(1): 115–135, July 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485714.

Anonymous:1985:Ea

[499] Anonymous. Erratum. Synthese, 64(2):ii, August 1985. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00486035;   http://link.
springer.com/content/pdf/10.1007/BF00486035.pdf.

Lycan:1985:EV

[500] William G. Lycan. Epistemic value. Synthese, 64(2):137–164, August 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00486036.

Sober:1985:PFP

[501] Elliott Sober. Panglossian functionalism and the philosophy of mind. Synthese, 64(2):165–193, August 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486037.

Bogen:1985:TEN

[502] James Bogen. Traditional epistemology and naturalistic replies to its skeptical critics. Synthese, 64(2):195–224, August 1985. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00486038.

Zemach:1985:N

[503] E. M. Zemach. Numbers. Synthese, 64(2):225–239, August 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00486039.

Wiesenthal:1985:VSP

[504] L. Wiesenthal. Visual space from the perspective of possible world semantics II. Synthese, 64(2):241–270, August 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00486040.

McGuinness:1985:P

[505] Brian McGuinness. Preface. Synthese, 64(3):271, September 1985. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485517.

Anonymous:1985:Ac

[506] Anonymous. Announcement. Synthese, 64(3):272, September 1985. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485518.

Geymonat:1985:DCS

[507] Ludovico Geymonat. Development and continuity in Schlick’s thought. Synthese, 64(3):273–282, September 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485519.

Haller:1985:PKM

[508] Rudolf Haller. Problems of knowledge in Moritz Schlick. Synthese, 64(3): 283–296, September 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485520.

Hung:1985:RA

[509] Tscha Hung. Remarks on affirmations. Synthese, 64(3):297–306, September 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485521.

vonKutschera:1985:MSS

[510] Franz von Kutschera. Moritz Schlick on self-evidence. Synthese, 64(3): 307–315, September 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485522.

Leinfellner:1985:RSP

[511] Werner Leinfellner. A reconstruction of Schlick’s psycho-sociological ethics. Synthese, 64(3):317–349, September 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485523.

McGuinness:1985:WVC

[512] Brian McGuinness. Wittgenstein and the Vienna circle. Synthese, 64(3): 351–358, September 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485524.

Melzer:1985:P

[513] Heinrich Melzer and Josef Schächter. On physicalism. Synthese, 64(3): 359–374, September 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485525.

Mulder:1985:VCA

[514] Henk Mulder. The Vienna Circle Archive and the literary remains of Moritz Schlick and Otto Neurath. Synthese, 64(3):375–387, September 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485526.

Quinton:1985:SBW

[515] Anthony Quinton. Schlick before Wittgenstein. Synthese, 64(3):389– 410, September 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485527.

Schleichert:1985:CUC

[516] Hubert Schleichert. On the concept of unity of consciousness. Synthese, 64 (3):411–420, September 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485528.

Richards:1985:P

[517] Barry Richards and Henri Wermus. Preface. Synthese, 65(1):1, October 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484884.

Kitchener:1985:GEH

[518] Richard F. Kitchener. Genetic epistemology, history of science and genetic psychology. Synthese, 65(1):3–31, October 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484885.

Richards:1985:CLR

[519] Barry Richards. Constructivism and logical reasoning. Synthese, 65(1): 33–64, October 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484886.

Haroutunian:1985:CJP

[520] Sophie Haroutunian. Can Jean Piaget explain the possibility of knowledge? Synthese, 65(1):65–86, October 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484887.

Matthews:1985:ICD

[521] Gareth B. Matthews. The idea of conceptual development in Piaget. Synthese, 65(1):87–97, October 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484888.

Buscaglia:1985:BJP

[522] Marino Buscaglia. La biologie de Jean Piaget: (1896–1980) — Cohérence et Marginalité. (French) [The biology of Jean Piaget: (1896–1980) — Coherence and Marginality]. Synthese, 65(1):99–120, October 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484889.

Voneche:1985:JPC

[523] Jacques Vonèche and Fernando Vidal. Jean Piaget and the child psychologist. Synthese, 65(1):121–138, October 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484890.

Kitchener:1985:BPW

[524] Richard F. Kitchener. Bibliography of philosophical work on Piaget. Synthese, 65(1):139–151, October 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484891.

Anonymous:1985:IUH

[525] Anonymous. International Union of the History and Philosophy of Science Division of Logic, Methodology and Philosophy of Science Bulletin No. 9. Synthese, 65(1):153–155, October 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484892.

Anonymous:1985:Eb

[526] Anonymous. Errata. Synthese, 65(1):156, October 1985. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484893;   http://link.
springer.com/content/pdf/10.1007/BF00484893.pdf.

Talja:1985:P

[527] Jari Talja and Douglas N. Walton. Preface. Synthese, 65(2):157, November 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869297.

Audi:1985:RR

[528] Robert Audi. Rationalization and rationality. Synthese, 65(2):159–184, November 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869298.

Segerberg:1985:R

[529] Krister Segerberg. Routines. Synthese, 65(2):185–210, November 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869299.

Walton:1985:PIA

[530] Douglas N. Walton. Pragmatic inferences about actions. Synthese, 65(2): 211–233, November 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869300.

Talja:1985:LO

[531] Jari Talja. On the logic of omissions. Synthese, 65(2):235–248, November 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869301.

Lehrer:1985:IIS

[532] Keith Lehrer and Carl Wagner. Intransitive indifference: The semiorder problem. Synthese, 65(2):249–256, November 1985. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869302.

Jennings:1985:CTN

[533] R. E. Jennings. Can there be a natural deontic logic? Synthese, 65(2): 257–273, November 1985. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869303.

Jones:1985:ISI

[534] Andrew J. I. Jones and Ingmar Pörn. Ideality, sub-ideality and deontic logic. Synthese, 65(2):275–290, November 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869304.

Aaqvist:1985:LSL

[535] Lennart Åqvist. On the logical syntax or linguistic deep structure of certain crime descriptions: Prolegomena to the doctrine of criminal intent. Synthese, 65(2):291–306, November 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869305.

Weinberger:1985:FRA

[536] Ota Weinberger. Freedom, range for action, and the ontology of norms. Synthese, 65(2):307–324, November 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869306.

Anonymous:1985:PR

[537] Anonymous. Press release. Synthese, 65(2):325, November 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869307.

Hausman:1985:EGD

[538] David B. Hausman. The explanation of goal-directed behavior. Synthese, 65(3):327–346, December 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869274.

Cramer:1985:EAN

[539] Jacqueline Cramer and Wolfgang van den Daele. Is ecology an ‘alternative’ natural science? Synthese, 65(3):347–375, December 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869275.

Clark:1985:QPE

[540] Austen Clark. Qualia and the psychophysiological explanation of color perception. Synthese, 65(3):377–405, December 1985. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869276.

Olding:1985:RNS

[541] A. Olding. Reductionism and natural selection. Synthese, 65(3):407– 410, December 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869277.

Alley:1985:OEM

[542] Thomas R. Alley. Organism-environment mutuality epistemics, and the concept of an ecological niche. Synthese, 65(3):411–444, December 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869278.

Edwards:1985:ADF

[543] David A. Edwards. Atomic discourse in The Feynman Lectures on Physics. Synthese, 65(3):445–480, December 1985. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869279.

Niiniluoto:1986:P

[544] Ilkka Niiniluoto. Preface. Synthese, 66(1):1, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413574.

vonWright:1986:TNC

[545] G. H. von Wright. Truth, negation, and contradiction. Synthese, 66(1): 3–14, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413575.

Back:1986:CIT

[546] R. J. R. Back. A computational interpretation of truth logic. Synthese, 66(1):15–34, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413576.

Bocharov:1986:BAS

[547] V. A. Bocharov. Boolean algebra and syllogism. Synthese, 66(1):35– 54, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413577.

Sidorenko:1986:STC

[548] E. A. Sidorenko. Semantic truth conditionals and relevant calculi. Synthese, 66(1):55–62, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413578.

Karpenko:1986:PSI

[549] A. S. Karpenko. Paraconsistent structure inside of many-valued logic. Synthese, 66(1):63–69, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413579.

Smirnov:1986:LRB

[550] V. A. Smirnov. Logical relations between theories. Synthese, 66(1):71– 87, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413580.

Jones:1986:OM

[551] Andrew J. I. Jones and Ingmar Pörn. ‘Ought’ and ‘must’. Synthese, 66 (1):89–93, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413581.

Aaqvist:1986:SRD

[552] Lennart Åqvist. Some results on dyadic deontic logic and the logic of preference. Synthese, 66(1):95–110, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413582.

Niiniluoto:1986:HIC

[553] Ilkka Niiniluoto. Hypothetical imperatives and conditional obligations. Synthese, 66(1):111–133, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413583.

Belzer:1986:RDP

[554] Marvin Belzer. Reasoning with defeasible principles. Synthese, 66(1):135– 158, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413584.

Mckinsey:1986:MA

[555] Michael Mckinsey. Mental anaphora. Synthese, 66(1):159–175, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413585.

Hyvonen:1986:ALI

[556] Eero Hyvönen. Applying a logical interpretation of semantic nets and graph grammars to natural language parsing and understanding. Synthese, 66(1):177–190, January 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413586.

Anonymous:1986:Aa

[557] Anonymous. Announcement. Synthese, 66(1):191–199, January 1986. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413587.

ONeill:1986:SBM

[558] John O’Neill. The specular body: Merleau–Ponty and Lacan on infant self and other. Synthese, 66(2):201–217, February 1986. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413644.

Sheets-Johnstone:1986:EFE

[559] Maxine Sheets-Johnstone. Existential fit and evolutionary continuities. Synthese, 66(2):219–248, February 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413645.

Rouse:1986:MPE

[560] Joseph Rouse. Merleau–Ponty and the existential conception of science. Synthese, 66(2):249–272, February 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413646.

Cunningham:1986:RRV

[561] Suzanne Cunningham. Representation: Rorty vs. Husserl. Synthese, 66 (2):273–289, February 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413647.

Harvey:1986:HPT

[562] Charles W. Harvey. Husserl and the problem of theoretical entities. Synthese, 66(2):291–309, February 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413648.

Pietersma:1986:HCE

[563] Henry Pietersma. Husserl’s concept of existence. Synthese, 66(2):311– 328, February 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413649.

Garrison:1986:HGP

[564] James W. Garrison. Husserl, Galileo, and the processes of idealization. Synthese, 66(2):329–338, February 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413650.

Levison:1986:MDM

[565] Arnold B. Levison. Metalinguistic dualism and the mark of the mental. Synthese, 66(3):339–359, March 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00414056.

Owens:1986:SNM

[566] Joseph Owens. Synonymy and the nonindividualistic model of the mental. Synthese, 66(3):361–382, March 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00414057.

Thurston:1986:SLA

[567] B. C. Thurston. On Sellars’ linguistic account of awareness. Synthese, 66(3):383–400, March 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414058.

Richards:1986:ICD

[568] Belinda Richards. Intentional content and demonstrative thought. Synthese, 66(3):401–404, March 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00414059.

Slezak:1986:ACS

[569] Peter Slezak. Actions, cognition and the self. Synthese, 66(3):405– 435, March 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414060.

Kirk:1986:MMG

[570] Robert Kirk. Mental machinery and Gödel. Synthese, 66(3):437– 452, March 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414061.

Yu:1986:CD

[571] Paul Yu and Gary Fuller. A critique of Dennett. Synthese, 66(3):453– 476, March 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414062.

Eldridge:1986:MIP

[572] Richard Eldridge. Metaphysics and the interpretation of persons: Davidson on thinking and conceptual schemes. Synthese, 66(3):477–503, March 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00414063.

Holman:1986:MM

[573] Emmett L. Holman. Maxwell and materialism. Synthese, 66(3):505– 514, March 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414064.

Burian:1986:I

[574] Richard M. Burian and Joseph C. Pitt. Introduction. Synthese, 67(1): 1–2, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485505.

Yolton:1986:THP

[575] John W. Yolton. Is there a history of philosophy? Some difficulties and suggestions. Synthese, 67(1):3–21, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485506.

Williams:1986:CJY

[576] W. H. Williams. Comment on John Yolton’s ‘is there a history of philosophy? Some difficulties and suggestions’. Synthese, 67(1):23–32, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485507.

Curley:1986:DD

[577] Edwin Curley. Dialogues with the dead. Synthese, 67(1):33–49, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485508.

Cohen:1986:DPD

[578] Lesley Cohen. Doing philosophy is doing its history. Synthese, 67(1): 51–55, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485509.

Turnbull:1986:WTR

[579] Robert G. Turnbull. On what there is: Representation and history. Synthese, 67(1):57–75, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485510.

Ariew:1986:DCG

[580] Roger Ariew. Descartes as critic of Galileo’s scientific methodology. Synthese, 67(1):77–90, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485511.

Garber:1986:LPR

[581] Daniel Garber. Learning from the past: Reflections on the role of history in the philosophy of science. Synthese, 67(1):91–114, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485512.

Laudan:1986:SPF

[582] Larry Laudan. Some problems facing intuitionist meta-methodologies. Synthese, 67(1):115–129, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485513.

Penelhum:1986:SBH

[583] Terence Penelhum. Sceptics, believers, and historical mistakes. Synthese, 67(1):131–146, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485514.

Stump:1986:PSF

[584] Eleonore Stump. Penelhum on skeptics and fideists. Synthese, 67(1): 147–154, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485515.

Anonymous:1986:Ab

[585] Anonymous. Announcement. Synthese, 67(1):155, April 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485516.

Ellett:1986:CPC

[586] Frederick S. Ellett, Jr. and David P. Ericson. Correlation, partial correlation, and causation. Synthese, 67(2):157–173, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00540066.

Ellett:1986:APC

[587] Frederick S. Ellett, Jr. and David P. Ericson. An analysis of probabilistic causation in dichotomous structures. Synthese, 67(2):175–193, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00540067.

Price:1986:ACD

[588] Huw Price. Against causal decision theory. Synthese, 67(2):195–212, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00540068.

Lee:1986:CCC

[589] Jig-Chuen Lee. Causal condition, causal asymmetry, and the counterfactual analysis of causation. Synthese, 67(2):213–223, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00540069.

Bishop:1986:ACC

[590] John Bishop. Is agent-causality a conceptual primitive? Synthese, 67(2): 225–247, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540070.

Ehring:1986:TTC

[591] Douglas Ehring. The transference theory of causation. Synthese, 67(2): 249–258, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540071.

Panda:1986:RSS

[592] Santosh C. Panda. On ranking sets of statements in terms of plausibility. Synthese, 67(2):259–271, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00540072.

Festa:1986:MDB

[593] Roberto Festa. A measure for the distance between an interval hypothesis and the truth. Synthese, 67(2):273–320, May 1986. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00540073.

Niiniluoto:1986:TBE

[594] Ilkka Niiniluoto. Truthlikeness and Bayesian estimation. Synthese, 67(2): 321–346, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540074.

Levi:1986:EEF

[595] Isaac Levi. Estimation and error free information. Synthese, 67(2): 347–360, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540075.

Fishburn:1986:OPD

[596] Peter C. Fishburn. Ordered preference differences without ordered preferences. Synthese, 67(2):361–368, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00540076.

Witmer:1986:OSS

[597] Peffrey A. Witmer and Murray K. Clayton. On objectivity and subjectivity in statistical inference: A response to Mayo. Synthese, 67(2):

369–379, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540077. See [348].

Anonymous:1986:CP

[598] Anonymous. Call for papers. Synthese, 67(2):380, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00540078.

Hilpinen:1986:IUH

[599] Risto Hilpinen. International Union of the History and Philosophy of Science Division of Logic, Methodology and Philosophy of Science Bulletin No. 10. Synthese, 67(2):381–382, May 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00540079.

Mundy:1986:EUR

[600] Brent Mundy. Embedding and uniqueness in relational theories of space. Synthese, 67(3):383–390, June 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485941.

Mundy:1986:GTM

[601] Brent Mundy. On the general theory of meaningful representation. Synthese, 67(3):391–437, June 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485942.

Adams:1986:CIA

[602] Ernest W. Adams. Continuity and idealizability of approximate generalizations. Synthese, 67(3):439–476, June 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485943.

deMuynck:1986:RBI

[603] Willem M. de Muynck and Gidi P. van Liempd. On the relation between indistinguishability of identical particles and (anti)symmetry of the wave function in quantum mechanics. Synthese, 67(3):477–496, June 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485944.

Mittelstaedt:1986:EAF

[604] Peter Mittelstaedt. Empiricism and apriorism in the foundations of quantum logic. Synthese, 67(3):497–525, June 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485945.

Tobar:1986:R

[605] José Felix Tobar. Review. Synthese, 67(3):527–530, June 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485946.

Humphreys:1986:CSS

[606] Paul Humphreys. Causation in the social sciences: An overview. Synthese, 68(1):1–12, July 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413963.

Anonymous:1986:EN

[607] Anonymous. Editor’s note. Synthese, 68(1):12, July 1986. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413964.

Blalock:1986:MCI

[608] Hubert M. Blalock, Jr. Multiple causation, indirect measurement and generalizability in the social sciences. Synthese, 68(1):13–36, July 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413965.

Glymour:1986:CMT

[609] Clark Glymour and Richard Scheines. Causal modeling with the TETRAD program. Synthese, 68(1):37–63, July 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413966.

Duncan:1986:PDI

[610] Otis Dudley Duncan. Probability, disposition, and the inconsistency of attitudes and behavior. Synthese, 68(1):65–98, July 1986. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413967.

Fetzer:1986:MIS

[611] James H. Fetzer. Methodological individualism: Singular causal systems and their population manifestations. Synthese, 68(1):99–128, July 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413968.

Suppes:1986:NMC

[612] Patrick Suppes. Non-Markovian causality in the social sciences with some theorems on transitivity. Synthese, 68(1):129–140, July 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413969.

Cook:1986:CAQ

[613] Thomas D. Cook and Donald T. Campbell. The causal assumptions of quasi-experimental practice. Synthese, 68(1):141–180, July 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413970.

Pettit:1986:PPD

[614] Philip Pettit. Preserving the prisoner’s dilemma. Synthese, 68(1):181– 184, July 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413971.

Makinson:1986:HGI

[615] David Makinson. How to give it up: A survey of some formal aspects of the logic of theory change. Synthese, 68(1):185–186, July 1986. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413972;   http:
//link.springer.com/content/pdf/10.1007/BF00413972.pdf.

Anonymous:1986:Ac

[616] Anonymous. Announcement. Synthese, 68(1):187, July 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413973.

Moser:1986:AEF

[617] Paul K. Moser. Ascending from empirical foundations. Synthese, 68(2): 189–203, August 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413830.

Keller:1986:FCJ

[618] James A. Keller. Foundationalism, circular justification, and the levels gambit. Synthese, 68(2):205–212, August 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413831.

Almeder:1986:FCR

[619] Robert Almeder. Fallibilism, coherence and realism. Synthese, 68(2): 213–223, August 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413832.

Siegel:1986:RTI

[620] Harvey Siegel. Relativism, truth, and incoherence. Synthese, 68(2):225– 259, August 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413833.

Garrison:1986:PIS

[621] James W. Garrison. The paradox of indoctrination: A solution. Synthese, 68(2):261–273, August 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413834.

Brown:1986:SCC

[622] Harold I. Brown. Sellars, concepts and conceptual change. Synthese, 68 (2):275–307, August 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413835.

Schlesinger:1986:PAP

[623] George Schlesinger. Present and absent properties. Synthese, 68(2):309– 331, August 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413836.

White:1986:CQ

[624] Stephen L. White. Curse of the qualia. Synthese, 68(2):333–368, August 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413837.

Shatz:1986:CEP

[625] David Shatz. Circularity and epistemic principles: A reply to James Keller. Synthese, 68(2):369–382, August 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413838.

Wahl:1986:BRT

[626] Russell Wahl. Bertrand Russell’s theory of judgment. Synthese, 68(3): 383–407, September 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413618.

Gert:1986:WPL

[627] Bernard Gert. Wittgenstein’s private language arguments. Synthese, 68 (3):409–439, September 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413619.

Roberts:1986:FGM

[628] Lawrence D. Roberts. The figure-ground model for the explanation of the determination of indexical reference. Synthese, 68(3):441–486, September 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413620.

Burns:1986:VC

[629] Linda Burns. Vagueness and coherence. Synthese, 68(3):487–513, September 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413621.

Indurkhya:1986:CST

[630] Bipin Indurkhya. Constrained semantic transference: A formal theory of metaphors. Synthese, 68(3):515–551, September 1986. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413622.

Sharrock:1986:MGS

[631] W. W. Sharrock and R. J. Anderson. Margaret Gilbert on the social nature of language. Synthese, 68(3):553–558, September 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413623.

Smith:1986:PRR

[632] Michael A. Smith. Peacocke on red and red’. Synthese, 68(3):559–576, September 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413624.

Peacocke:1986:RMS

[633] Christopher Peacocke. Reply to Michael Smith. Synthese, 68(3):577– 580, September 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413625.

Gauker:1986:PC

[634] Christopher Gauker. The principle of charity. Synthese, 69(1):1–25, October 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01988284.

Gibson:1986:QD

[635] Roger F. Gibson. Quine’s dilemma. Synthese, 69(1):27–39, October 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF01988285.

Harris:1986:LLG

[636] James F. Harris. Language, language games and ostensive definition. Synthese, 69(1):41–49, October 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01988286.

Schatzki:1986:RMU

[637] Theodore R. Schatzki. The rationalization of meaning and understanding: Davidson and Habermas. Synthese, 69(1):51–79, October 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01988287.

Siderits:1986:SRD

[638] Mark Siderits. The sense-reference distinction in Indian philosophy of language. Synthese, 69(1):81–106, October 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF01988288.

Stampe:1986:VCA

[639] Dennis Stampe. Verificationism and a causal account of meaning. Synthese, 69(1):107–137, October 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF01988289.

Anonymous:1986:LLC

[640] Anonymous. Logic and linguistics conference. Synthese, 69(1):138, October 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF01988290.

Anonymous:1986:LA

[641] Anonymous. Lakatos Award. Synthese, 69(1):139, October 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF01988291.

Laudan:1986:SCP

[642] Larry Laudan, Arthur Donovan, Rachel Laudan, Peter Barker, Harold Brown, Jarrett Leplin, Paul Thagard, and Steve Wykstra. Scientific change: Philosophical models and historical research. Synthese, 69(2): 141–223, November 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413981.

Doppelt:1986:RRM

[643] Gerald Doppelt. Relativism and the reticulational model of scientific rationality. Synthese, 69(2):225–252, November 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413982.

Nickles:1986:RUH

[644] Thomas Nickles. Remarks on the use of history as evidence. Synthese, 69 (2):253–266, November 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413983.

Hall:1986:PEP

[645] Peter Joseph Hall. The Pauli Exclusion Principle and the foundations of chemistry. Synthese, 69(3):267–272, December 1986. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413974.

Vermeeren:1986:CEC

[646] Hans P. W. Vermeeren. Controversies and existence claims in chemistry: The theory of resonance. Synthese, 69(3):273–290, December 1986. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413975.

VanBrakel:1986:CSP

[647] J. Van Brakel. The chemistry of substances and the philosophy of mass terms. Synthese, 69(3):291–324, December 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413976.

Halpin:1986:SCB

[648] John F. Halpin. Stalnaker’s conditional and Bell’s problem. Synthese, 69 (3):325–340, December 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413977.

Tait:1986:TPP

[649] W. W. Tait. Truth and proof: The Platonism of mathematics. Synthese, 69 (3):341–370, December 1986. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413978.

Friedman:1986:HFF

[650] Joel Friedman. How the finite follows from the infinite in Spinoza’s metaphysical system. Synthese, 69(3):371–407, December 1986. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413979.

Bell:1986:ALM

[651] J. L. Bell. From absolute to local mathematics. Synthese, 69(3):409– 426, December 1986. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413980.

Stabler:1987:KFA

[652] Edward P. Stabler, Jr. Kripke on functionalism and automata. Synthese, 70(1):1–22, January 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414025.

Lloyd:1987:MRB

[653] Dan Lloyd. Mental representation from the bottom up. Synthese, 70 (1):23–78, January 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414026.

Demopoulos:1987:SFD

[654] William Demopoulos. On some fundamental distinctions of computationalism. Synthese, 70(1):79–96, January 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414027.

Pylyshyn:1987:WM

[655] Zenon W. Pylyshyn. What’s in a mind? Synthese, 70(1):97–122, January 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00414028.

Cussins:1987:VP

[656] Adrian Cussins. Varieties of psychologism. Synthese, 70(1):123–154, January 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414029.

Anonymous:1987:Aa

[657] Anonymous. Announcements. Synthese, 70(1):155–156, January 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00414030.

Loewer:1987:P

[658] Barry Loewer. Preface. Synthese, 70(2):157, February 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413932.

Foley:1987:DIT

[659] Richard Foley. Dretske’s “information-theoretic” account of knowledge. Synthese, 70(2):159–184, February 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413933.

Savitt:1987:AIC

[660] Steven F. Savitt. Absolute informational content. Synthese, 70(2):185– 190, February 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413934.

Grandy:1987:IBE

[661] Richard E. Grandy. Information-based epistemology, ecological epistemology and epistemology naturalized. Synthese, 70(2):191–203, February 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413935.

Bogdan:1987:MCI

[662] Radu J. Bogdan. Mind, content and information. Synthese, 70(2):205– 227, February 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413936.

Yourgrau:1987:IRC

[663] Palle Yourgrau. Information retrieval and cognitive accessibility. Synthese, 70(2):229–246, February 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413937.

Sayre:1987:CSP

[664] Ken Sayre. Cognitive science and the problem of semantic content. Synthese, 70(2):247–269, February 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413938.

McLaughlin:1987:WWC

[665] Brian P. McLaughlin. What is wrong with correlational psychosemantics. Synthese, 70(2):271–286, February 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413939.

Loewer:1987:II

[666] Barry Loewer. From information to intentionality. Synthese, 70(2):287– 317, February 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413940.

Anonymous:1987:Ab

[667] Anonymous. Announcements. Synthese, 70(2):318, February 1987. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413941.

Osherson:1987:SLC

[668] D. N. Osherson, M. Stob, and S. Weinstein. Social learning and collective choice. Synthese, 70(3):319–347, March 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414154.

Maloney:1987:RS

[669] J. Christopher Maloney. The right stuff. Synthese, 70(3):349–372, March 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00414155.

Mundy:1987:FRP

[670] Brent Mundy. Faithful representation, physical extensive measurement theory and Archimedean axioms. Synthese, 70(3):373–400, March 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00414156.

Russow:1987:SFC

[671] Lilly-Marlene Russow. Stich on the foundations of cognitive psychology. Synthese, 70(3):401–413, March 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00414157.

Talbott:1987:SNS

[672] W. J. Talbott. Standard and non-standard Newcomb problems. Synthese, 70(3):415–458, March 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414158.

Ruse:1987:R

[673] Michael Ruse. Review. Synthese, 70(3):459–462, March 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00414159.

Anonymous:1987:NPRa

[674] Anonymous. Note from the publisher regarding volume contents list. Synthese, 70(3):463, March 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414160.

Bacon:1987:MTC

[675] John Bacon. A model-theoretic criterion of ontology. Synthese, 71(1): 1–18, April 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486433.

Cover:1987:CPC

[676] J. A. Cover. Causal priority and causal conditionship. Synthese, 71(1): 19–36, April 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486434.

Brady:1987:JMK

[677] Michael E. Brady. J. M. Keynes’ ‘theory of evidential weight’: Its relation to information processing theory and application in the general theory. Synthese, 71(1):37–59, April 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00486435.

Pollock:1987:EN

[678] John L. Pollock. Epistemic norms. Synthese, 71(1):61–95, April 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486436.

Taylor:1987:BIS

[679] Kenneth A. Taylor. Belief, information and semantic content: A naturalist’s lament. Synthese, 71(1):97–124, April 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00486437.

Scriba:1987:A

[680] C. J. Scriba. Announcement. Synthese, 71(1):125–126, April 1987. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486438.

Anonymous:1987:FLP

[681] Anonymous. First lecture practice based on theory as the object of science. Synthese, 71(2):127–139, May 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00486186.

Anonymous:1987:SLP

[682] Anonymous. Second lecture political practice and ethico-political theory. Synthese, 71(2):140–155, May 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486187.

Anonymous:1987:TLS

[683] Anonymous. Third lecture social technology and statistics. Synthese, 71(2):156–172, May 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486188.

Anonymous:1987:FLT

[684] Anonymous. Fourth lecture technological practice and physical/ technological theories. Synthese, 71(2):173–186, May 1987. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00486189.

Anonymous:1987:FLG

[685] Anonymous. Fifth lecture geometry and kinematics. Synthese, 71(2): 187–203, May 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486190.

Lorenzen:1987:SLC

[686] P. Lorenzen. Sixth lecture collision mechanics and dynamics. Synthese, 71(2):204–218, May 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486191.

Anonymous:1987:Ac

[687] Anonymous. Announcement. Synthese, 71(2):219, May 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00486192.

Laudan:1987:RNR

[688] Larry Laudan. Relativism, naturalism and reticulation. Synthese, 71(3): 221–234, June 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485630.

Blachowicz:1987:DC

[689] James Blachowicz. Discovery as correction. Synthese, 71(3):235–321, June 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485631.

Goldman:1987:FPL

[690] Alan H. Goldman. The force of precedent in legal, moral, and empirical reasoning. Synthese, 71(3):323–346, June 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485632.

Braaten:1987:RCP

[691] Jane Braaten. Rational consensual procedure: Argumentation or weighted averaging? Synthese, 71(3):347–354, June 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485633.

Berger:1987:SVS

[692] George Berger. On the structure of visual sentience. Synthese, 71(3): 355–370, June 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485634.

Thomas:1987:T

[693] Laurence Thomas. A tribute. Synthese, 72(1):1–4, July 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413812.

Gauthier:1987:RM

[694] David Gauthier. Reason to be moral? Synthese, 72(1):5–27, July 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413813.

Rescher:1987:RMO

[695] Nicholas Rescher. Rationality and moral obligation. Synthese, 72(1): 29–43, July 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413814.

Sterba:1987:JMR

[696] James P. Sterba. Justifying morality: The right and the wrong ways. Synthese, 72(1):45–69, July 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413815.

Darwall:1987:AM

[697] Stephen L. Darwall. Abolishing morality. Synthese, 72(1):71–89, July 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413816.

Richards:1987:MR

[698] David A. J. Richards. Moral rationality. Synthese, 72(1):91–101, July 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413817.

Thomson:1987:VA

[699] Judith Jarvis Thomson. Verbs of action. Synthese, 72(1):103–122, July 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413818.

Schneewind:1987:PPH

[700] J. B. Schneewind. Pufendorf’s place in the history of ethics. Synthese, 72(1):123–155, July 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413819.

Baier:1987:CL

[701] Annette C. Baier. Commodious living. Synthese, 72(2):157–185, August 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413637.

Cooper:1987:CHR

[702] John M. Cooper. Contemplation and happiness: A reconsideration. Synthese, 72(2):187–216, August 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413638.

Thomas:1987:F

[703] Laurence Thomas. Friendship. Synthese, 72(2):217–236, August 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413639.

Benditt:1987:LM

[704] Theodore M. Benditt. Liberal morality. Synthese, 72(2):237–247, August 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413640.

Feinberg:1987:SUD

[705] Joel Feinberg. Some unswept debris from the Hart–Devlin debate. Synthese, 72(2):249–275, August 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413641.

Stocker:1987:SPC

[706] Michael Stocker. Some problems with counter-examples in ethics. Synthese, 72(2):277–289, August 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413642.

Anonymous:1987:Ad

[707] Anonymous. Announcements. Synthese, 72(2):291–292, August 1987. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413643.

Bechtel:1987:PCC

[708] William Bechtel. Psycholinguistics as a case of cross-disciplinary research: Symposium introduction. Synthese, 72(3):293–311, September 1987. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413748.

Blumenthal:1987:EP

[709] Arthur L. Blumenthal. The emergence of psycholinguistics. Synthese, 72 (3):313–323, September 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413749.

Reber:1987:RSR

[710] Arthur S. Reber. The rise and (surprisingly rapid) fall of psycholinguistics. Synthese, 72(3):325–339, September 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413750.

McCauley:1987:HSM

[711] Robert N. McCauley. The not so happy story of the marriage of linguistics and psychology or why linguistics has discouraged psychology’s recent advances. Synthese, 72(3):341–353, September 1987. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413751.

Abrahamsen:1987:BBV

[712] Adele A. Abrahamsen. Bridging boundaries versus breaking boundaries: Psycholinguistics in perspective. Synthese, 72(3):355–388, September 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413752.

Anonymous:1987:NPRb

[713] Anonymous. Note from the publisher regarding volume contents list. Synthese, 72(3):389, September 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413753.

Schmitt:1987:I

[714] Frederick F. Schmitt. Introduction. Synthese, 73(1):1–2, October 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485439.

Cohen:1987:KCS

[715] Stewart Cohen. Knowledge, context, and social standards. Synthese, 73(1):3–26, October 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485440.

Kornblith:1987:SSF

[716] Hilary Kornblith. Some social features of cognition. Synthese, 73(1): 27–41, October 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485441.

Schmitt:1987:JSA

[717] Frederick F. Schmitt. Justification, sociality, and autonomy. Synthese, 73(1):43–85, October 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485442.

Lehrer:1987:PSK

[718] Keith Lehrer. Personal and social knowledge. Synthese, 73(1):87– 107, October 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485443.

Goldman:1987:FSE

[719] Alvin I. Goldman. Foundations of social epistemics. Synthese, 73(1):109– 144, October 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485444.

Fuller:1987:RWK

[720] Steve Fuller. On regulating what is known: A way to social epistemology. Synthese, 73(1):145–183, October 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485445.

Gilbert:1987:MCB

[721] Margaret Gilbert. Modelling collective belief. Synthese, 73(1):185– 204, October 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485446.

Carruthers:1987:CP

[722] Peter Carruthers. Conceptual pragmatism. Synthese, 73(2):205–224, November 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484740.

Henderson:1987:PCP

[723] David K. Henderson. The principle of charity and the problem of irrationality (translation and the problem of irrationality). Synthese, 73(2): 225–252, November 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484741.

Werhane:1987:SPK

[724] Patricia H. Werhane. Some paradoxes in Kripke’s interpretation of Wittgenstein. Synthese, 73(2):253–273, November 1987. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484742.

Haack:1987:R

[725] Susan Haack. Realism. Synthese, 73(2):275–299, November 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484743.

Gemes:1987:WIN

[726] Ken Gemes. The world in itself: Neither uniform nor physical. Synthese, 73(2):301–318, November 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484744.

Ehring:1987:CR

[727] Douglas Ehring. Causal relata. Synthese, 73(2):319–328, November 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484745.

Milne:1987:PP

[728] Peter Milne. Physical probabilities. Synthese, 73(2):329–359, November 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484746.

Bigelow:1987:AIP

[729] John Bigelow and Robert Pargetter. An analysis of indefinite probability statements. Synthese, 73(2):361–370, November 1987. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484747.

Redei:1987:RHV

[730] Miklós Rédei. Reformulation of the hidden variable problem using entropic measure of uncertainty. Synthese, 73(2):371–379, November 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484748.

Dieks:1987:GUF

[731] Dennis Dieks. Gravitation as a universal force. Synthese, 73(2):381– 397, November 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484749.

Simberloff:1987:R

[732] Daniel Simberloff. Review. Synthese, 73(2):399–405, November 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484750.

Martin-Lof:1987:TPE

[733] Per Martin-Löf. Truth of a proposition, evidence of a judgement, validity of a proof. Synthese, 73(3):407–420, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:

//link.springer.com/article/10.1007/BF00484985.

Casari:1987:CL

[734] Ettore Casari. Comparative logics. Synthese, 73(3):421–449, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484986.

vanBenthem:1987:MII

[735] Johan van Benthem. Meaning: Interpretation and inference. Synthese, 73 (3):451–470, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484987.

Prawitz:1987:SRV

[736] Dag Prawitz. Some remarks on verificationistic theories of meaning. Synthese, 73(3):471–477, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484988.

Chiara:1987:AIS

[737] Maria Luisa Dalla Chiara. An approach to intensional semantics. Synthese, 73(3):479–496, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484989.

Hintikka:1987:LUS

[738] Jaakko Hintikka. Language understanding and strategic meaning. Synthese, 73(3):497–529, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484990.

Jacob:1987:TPH

[739] Pierre Jacob. Is there a path half-way between realism and verificationism? Synthese, 73(3):531–547, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484991.

Eco:1987:MD

[740] Umberto Eco. Meaning and denotation. Synthese, 73(3):549–568, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484992.

Guenthner:1987:LMD

[741] Franz Guenthner. Linguistic meaning in discourse representation theory. Synthese, 73(3):569–598, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484993.

Bencivenga:1987:NPM

[742] Ermanno Bencivenga. A new paradigm of meaning. Synthese, 73(3): 599–621, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484994.

Bonomi:1987:RS

[743] Andrea Bonomi. Relevant situations. Synthese, 73(3):623–636, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484995.

Santambrogio:1987:GIO

[744] Marco Santambrogio. Generic and intensional objects. Synthese, 73(3): 637–663, December 1987. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484996.

Anonymous:1988:PAa

[745] Anonymous. Publisher’s announcement. Synthese, 74(1):1, January 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869614.

H:1988:I

[746] J. H. Introduction. Synthese, 74(1):3–4, January 1988. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869615.

Meyer:1988:RQT

[747] Michel Meyer. The revival of questioning in the twentieth century. Synthese, 74(1):5–18, January 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869616.

Kleiner:1988:ELS

[748] Scott A. Kleiner. Erotetic logic and scientific inquiry. Synthese, 74(1): 19–46, January 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869617.

Bromberger:1988:RI

[749] Sylvain Bromberger. Rational ignorance. Synthese, 74(1):47–64, January 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869618.

Zytkow:1988:NSD

[750] Jan M. Zytkow and Herbert A. Simon. Normative systems of discovery and logic of search. Synthese, 74(1):65–90, January 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869619.

Kevelson:1988:HWW

[751] Roberta Kevelson. How’s of why’s and why’s of how’s: Relation of method and cause in inquiry. Synthese, 74(1):91–106, January 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869620.

Loparic:1988:SPK

[752] Zelko Loparic. System-problems in Kant. Synthese, 74(1):107–140, January 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869621.

Anonymous:1988:Aa

[753] Anonymous. Announcement. Synthese, 74(1):141, January 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869622.

Jh:1988:I

[754] Jh. Introduction. Synthese, 74(2):143–144, February 1988. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869549.

Garrison:1988:HLN

[755] James W. Garrison. Hintikka, Laudan and Newton: An interrogative model of scientific inquiry. Synthese, 74(2):145–171, February 1988. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869550.

Hintikka:1988:WLE

[756] Jaakko Hintikka. What is the logic of experimental inquiry? Synthese, 74 (2):173–190, February 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869551.

Koura:1988:AWQ

[757] Antti Koura. An approach to why-questions. Synthese, 74(2):191– 206, February 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869552.

Hand:1988:GTS

[758] Michael Hand. Game-theoretical semantics, Montague semantics, and questions. Synthese, 74(2):207–222, February 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869553.

Carlson:1988:QHS

[759] Lauri Carlson. Quantified Hintikka-style epistemic logic. Synthese, 74 (2):223–262, February 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869554.

Luper-Foy:1988:P

[760] Steven Luper-Foy. Preface. Synthese, 74(3):263, March 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869629.

Alston:1988:IE

[761] William P. Alston. An internalist externalism. Synthese, 74(3):265– 283, March 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869630.

Chisholm:1988:IIJ

[762] Roderick M. Chisholm. The indispensability of internal justification. Synthese, 74(3):285–296, March 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869631.

Elgin:1988:EES

[763] Catherine Z. Elgin. The epistemic efficacy of stupidity. Synthese, 74 (3):297–311, March 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869632.

Kornblith:1988:HIC

[764] Hilary Kornblith. How internal can you get? Synthese, 74(3):313– 327, March 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869633.

Lehrer:1988:MUJ

[765] Keith Lehrer. Metaknowledge: Undefeated justification. Synthese, 74 (3):329–347, March 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869634.

Luper-Foy:1988:KI

[766] Steven Luper-Foy. The knower, inside and out. Synthese, 74(3):349– 367, March 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869635.

Pollock:1988:IDR

[767] John L. Pollock. Interest-driven reasoning. Synthese, 74(3):369–390, March 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869636.

Stich:1988:REA

[768] Stephen Stich. Reflective equilibrium, analytic epistemology and the problem of cognitive diversity. Synthese, 74(3):391–413, March 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869637.

Sosa:1988:MAB

[769] Ernest Sosa. Methodology and apt belief. Synthese, 74(3):415–426, March 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869638.

Mundy:1988:EMR

[770] Brent Mundy. Extensive measurement and ratio functions. Synthese, 75 (1):1–23, April 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873272.

Madan:1988:DTC

[771] Dilip B. Madan and J. C. Owings. Decision theory with complex uncertainties. Synthese, 75(1):25–44, April 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00873273.

Sorensen:1988:VMB

[772] Roy A. Sorensen. Vagueness, measurement, and blurriness. Synthese, 75(1):45–82, April 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873274.

Vinci:1988:OCI

[773] Thomas C. Vinci. Objective chance, indicative conditionals and decision theory; or, how you can be smart, rich and keep on smoking. Synthese, 75(1):83–105, April 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873275.

Sobel:1988:DCR

[774] Jordan Howard Sobel. Defenses and conservative revisions of evidential decision theories: Metatickles and ratificationism. Synthese, 75(1): 107–131, April 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873276.

Lennon:1988:IN

[775] Thomas M. Lennon and Robert E. Butts. Introductory note. Synthese, 75(2):133, May 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00872974.

Hacking:1988:LLL

[776] Ian Hacking. Locke, Leibniz, language and Hans Aarsleff. Synthese, 75(2): 135–153, May 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00872975.

McRae:1988:LLL

[777] Robert McRae. Locke and Leibniz on linguistic particles. Synthese, 75(2): 155–161, May 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00872976.

Duchesneau:1988:LCF

[778] François Duchesneau. Leibniz on the classificatory function of language. Synthese, 75(2):163–181, May 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00872977.

Burkhardt:1988:MLT

[779] Hans Burkhardt. Modalities in language, thought and reality in Leibniz, Descartes and Crusius. Synthese, 75(2):183–215, May 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00872978.

Ross:1988:HDR

[780] George Macdonald Ross. Hobbes and Descartes on the relation between language and consciousness. Synthese, 75(2):217–229, May 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00872979.

Lennon:1988:BI

[781] Thomas M. Lennon. Berkeley and the ineffable. Synthese, 75(2):231– 250, May 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00872980.

Butts:1988:GRH

[782] Robert E. Butts. The grammar of reason: Hamann’s challenge to Kant. Synthese, 75(2):251–283, May 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00872981.

Bell:1988:I

[783] J. L. Bell. Infinitesimals. Synthese, 75(3):285–315, June 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869403.

Fuhrmann:1988:PFL

[784] Gy. Fuhrmann. “Prototypes” and “fuzziness” in the logic of concepts. Synthese, 75(3):317–347, June 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869404.

Fuhrmann:1988:FCC

[785] Gy. Fuhrmann. Fuzziness of concepts and concepts of fuzziness. Synthese, 75(3):349–372, June 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869405.

Tieszen:1988:PMK

[786] Richard Tieszen. Phenomenology and mathematical knowledge. Synthese, 75(3):373–403, June 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869406.

Williamson:1988:BSC

[787] Timothy Williamson. Bivalence and subjunctive conditionals. Synthese, 75(3):405–421, June 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869407.

Malin:1988:R

[788] Shimon Malin. Review. Synthese, 75(3):423–428, June 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869408.

Brady:1988:JMK

[789] Michael Emmett Brady. J. M. Keynes’s position on the general applicability of mathematical, logical and statistical methods in economics and social science. Synthese, 76(1):1–24, July 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869639.

Chandler:1988:MVB

[790] Marthe Chandler. Models of voting behavior in survey research. Synthese, 76(1):25–48, July 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869640.

Levi:1988:ICR

[791] Isaac Levi. Iteration of conditionals and the Ramsey test. Synthese, 76 (1):49–81, July 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869641.

Kelley:1988:RDC

[792] Leigh B. Kelley. Reflections on deliberative coherence. Synthese, 76(1): 83–121, July 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869642.

Pettit:1988:PDU

[793] Philip Pettit. The prisoner’s dilemma is an unexploitable Newcomb problem. Synthese, 76(1):123–134, July 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869643.

Bicchieri:1988:SBC

[794] Cristina Bicchieri. Strategic behavior and counterfactuals. Synthese, 76 (1):135–169, July 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869644.

White:1988:UBR

[795] Michael J. White. The unimportance of being random. Synthese, 76(1): 171–178, July 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869645.

Kyburg:1988:R

[796] Henry E. Kyburg, Jr. Review. Synthese, 76(1):179–182, July 1988. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869646.

Dacey:1988:GEP

[797] Raymond Dacey. Guest editor’s preface. Synthese, 76(2):183–184, August 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869587.

Brams:1988:NSG

[798] Steven J. Brams and D. Marc Kilgour. National security games. Synthese, 76(2):185–200, August 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869588.

Hirshleifer:1988:ACC

[799] Jack Hirshleifer. The analytics of continuing conflict. Synthese, 76(2): 201–233, August 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869589.

Intriligator:1988:PPM

[800] Michael D. Intriligator and Dagobert L. Brito. A predator–prey model of guerrilla warfare. Synthese, 76(2):235–244, August 1988. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869590.

Sandler:1988:CDA

[801] Todd Sandler and Harvey E. Lapan. The calculus of dissent: An analysis of terrorists’ choice of targets. Synthese, 76(2):245–261, August 1988.

CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869591.

deMesquita:1988:ARO

[802] Bruce Bueno de Mesquita and David Lalman. Arms races and the opportunity for peace. Synthese, 76(2):263–283, August 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869592.

Cioffi-Revilla:1988:PWC

[803] Claudio Cioffi-Revilla and Raymond Dacey. The probability of war in the n-crises problem: Modeling new alternatives to Wright’s solution. Synthese, 76(2):285–305, August 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869593.

Zinnes:1988:WPI

[804] Dina A. Zinnes and Robert G. Muncaster. The war propensity of international systems. Synthese, 76(2):307–331, August 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:

//link.springer.com/article/10.1007/BF00869594.

Burks:1988:TLM

[805] Arthur W. Burks. Teleology and logical mechanism. Synthese, 76(3): 333–370, September 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869605.

Ehring:1988:CAC

[806] Douglas Ehring. Causal asymmetry and causal relata: Reply to Lee. Synthese, 76(3):371–375, September 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869606.

Ranta:1988:PGT

[807] Aarne Ranta. Propositions as games as types. Synthese, 76(3):377–395, September 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869607.

Sanford:1988:CTO

[808] David H. Sanford. Can there be one-way causal conditionship? Synthese, 76(3):397–408, September 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869608.

Ellis:1988:IR

[809] Brian Ellis. Internal realism. Synthese, 76(3):409–434, September 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869609.

Villela-Minnerly:1988:TSC

[810] Lucia Villela-Minnerly. A temporary suspension of certainties. Synthese, 76(3):435–440, September 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869610.

Shrader-Frechette:1988:R

[811] Kristin Shrader-Frechette. Review. Synthese, 76(3):441–446, September 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869611.

Almeder:1988:R

[812] Robert Almeder. Review. Synthese, 76(3):447–451, September 1988. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869612.

Anonymous:1988:IUH

[813] Anonymous. International Union of the History and Philosophy of Science Division of Logic, Methodology and Philosophy of Science (IUHPS/ DLMPS) Bulletin No. 12. Synthese, 76(3):453–473, September 1988. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869613.

Hintikka:1988:DMT

[814] Jaakko Hintikka. On the development of the model-theoretic viewpoint in logical theory. Synthese, 77(1):1–36, October 1988. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869545.

Cocchiarella:1988:PVM

[815] Nino B. Cocchiarella. Predication versus membership in the distinction between logic as language and logic as calculus. Synthese, 77(1): 37–72, October 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869546.

Haaparanta:1988:AML

[816] Leila Haaparanta. Analysis as the method of logical discovery: Some remarks on Frege and Husserl. Synthese, 77(1):73–97, October 1988. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869547.

Kusch:1988:HHM

[817] Martin Kusch. Husserl and Heidegger on meaning. Synthese, 77(1):99– 127, October 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869548.

Anonymous:1988:PAb

[818] Anonymous. Publisher’s announcement. Synthese, 77(2):iii, November 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869431.

Kalinowski:1988:PGP

[819] M. W. Kalinowski. The program of geometrization of physics: Some philosophical remarks. Synthese, 77(2):129–138, November 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869432.

Klotz:1988:NQM

[820] A. H. Klotz. On the nature of quantum mechanics. Synthese, 77(2): 139–193, November 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869433.

Albert:1988:IMW

[821] David Albert and Barry Loewer. Interpreting the many worlds interpretation. Synthese, 77(2):195–213, November 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869434.

Mormann:1988:SRC

[822] Thomas Mormann. Structuralist reduction concepts as structurepreserving maps. Synthese, 77(2):215–250, November 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:

//link.springer.com/article/10.1007/BF00869435.

Kincaid:1988:SE

[823] Harold Kincaid. Supervenience and explanation. Synthese, 77(2):251– 281, November 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869436.

Anonymous:1988:Ab

[824] Anonymous. Announcement. Synthese, 77(2):283, November 1988. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869437.

Anonymous:1988:CP

[825] Anonymous. Call for papers. Synthese, 77(2):284, November 1988. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869438.

dePierris:1988:FKP

[826] Graciela de Pierris. Frege and Kant on a priori knowledge. Synthese, 77 (3):285–319, December 1988. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869289.

Pereboom:1988:KI

[827] Derk Pereboom. Kant on intentionality. Synthese, 77(3):321–352, December 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869290.

Baldner:1988:CTT

[828] Kent Baldner. Causality and things in themselves. Synthese, 77(3):353– 373, December 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869291.

Krausser:1988:AAD

[829] Peter Krausser. On the antinomies and the appendix to the dialectic in Kant’s critique and philosophy of science. Synthese, 77(3):375– 401, December 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869292.

Wehrle:1988:R

[830] Walter Wehrle. Review. Synthese, 77(3):403–407, December 1988. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869293.

Nadler:1988:R

[831] Steven M. Nadler. Review. Synthese, 77(3):409–413, December 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869294.

Harvey:1988:R

[832] Charles W. Harvey. Review. Synthese, 77(3):415–425, December 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869295.

Anonymous:1988:Ac

[833] Anonymous. Announcement. Synthese, 77(3):426–427, December 1988. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869296.

Cushing:1989:JSS

[834] James T. Cushing. The justification and selection of scientific theories. Synthese, 78(1):1–24, January 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869679.

Mcallister:1989:TBS

[835] James W. Mcallister. Truth and beauty in scientific reason. Synthese, 78(1):25–51, January 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869680.

Rantala:1989:WFW

[836] Veikko Rantala and Liselotte Wiesenthal. The worlds of fiction and the worlds of science: A comparative study. Synthese, 78(1):53–86, January 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869681.

Stillwfii:1989:PAR

[837] Shelley I. Stillwfii. Plantinga and anti-realism. Synthese, 78(1):87– 115, January 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869682.

Kaelin:1989:R

[838] E. F. Kaelin. Review. Synthese, 78(1):117–125, January 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869683.

Anonymous:1989:Aa

[839] Anonymous. Announcement. Synthese, 78(1):I–II, January 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869684.

Doepke:1989:SI

[840] Frederick Doepke. The step to individuation. Synthese, 78(2):129– 140, February 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869369.

French:1989:WPI

[841] Steven French. Why the principle of the identity of indiscernibles is not contingently true either. Synthese, 78(2):141–166, February 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869370.

Smith:1989:MUI

[842] Quentin Smith. The multiple uses of indexicals. Synthese, 78(2):167– 191, February 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869371.

Strawson:1989:RR

[843] Galen Strawson. Red and ‘red’. Synthese, 78(2):193–232, February 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869372.

Anonymous:1989:Ab

[844] Anonymous. Announcement. Synthese, 78(2):I–II, February 1989. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869373.

Malpas:1989:INL

[845] J. E. Malpas. The intertranslatability of natural languages. Synthese, 78 (3):233–264, March 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869350.

Wrigley:1989:OWV

[846] Michael Wrigley. The origins of Wittgenstein’s verificationism. Synthese, 78(3):265–290, March 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869351.

Sievert:1989:ALW

[847] Don Sievert. Another look at Wittgenstein on color exclusion. Synthese, 78(3):291–318, March 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869352.

Klagge:1989:WN

[848] James C. Klagge. Wittgenstein and neuroscience. Synthese, 78(3):319– 343, March 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869353.

Addis:1989:R

[849] Laird Addis. Review. Synthese, 78(3):345–356, March 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869354.

Wilkins:1989:R

[850] Burleigh T. Wilkins. Review. Synthese, 78(3):357–358, March 1989. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869355.

Anonymous:1989:Ac

[851] Anonymous. Announcement. Synthese, 78(3):359–360, March 1989. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869356.

Sundholm:1989:CGQ

[852] Göran Sundholm. Constructive generalized quantifiers. Synthese, 79(1): 1–12, April 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873254.

Shapiro:1989:LOM

[853] Stewart Shapiro. Logic, ontology, mathematical practice. Synthese, 79 (1):13–50, April 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873255.

Stenlund:1989:CLS

[854] Sören Stenlund. On the concept of language in some recent theories of meaning. Synthese, 79(1):51–98, April 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00873256.

Mackenzie:1989:RL

[855] Jim Mackenzie. Reasoning and logic. Synthese, 79(1):99–117, April 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00873257.

Dietrich:1989:SCP

[856] Eric Dietrich. Semantics and the computational paradigm in cognitive psychology. Synthese, 79(1):119–141, April 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00873258.

Kvanvig:1989:CV

[857] Jonathan L. Kvanvig. Conservatism and its virtues. Synthese, 79(1): 143–163, April 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873259.

Bruce:1989:R

[858] Darryl Bruce. Review. Synthese, 79(1):165–169, April 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00873260.

Forbes:1989:R

[859] Graeme Forbes. Review. Synthese, 79(1):171–189, April 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00873261.

Ryckman:1989:BSE

[860] Thomas C. Ryckman. On believing, saying and expressing. Synthese, 79(2):191–200, May 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869623.

Warmbrod:1989:BSH

[861] Ken Warmbrod. Beliefs and sentences in the head. Synthese, 79(2): 201–230, May 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869624.

DAgostino:1989:AEC

[862] Fred D’Agostino. Adjudication as an epistemological concept. Synthese, 79(2):231–256, May 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869625.

Davis:1989:TFC

[863] Wayne A. Davis and John W. Bender. Technical flaws in the coherence theory. Synthese, 79(2):257–278, May 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869626.

Forrest:1989:PRI

[864] Peter Forrest. The problem of representing incompletely ordered doxastic systems. Synthese, 79(2):279–303, May 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869627.

Shelton:1989:STK

[865] Jim Shelton. Schlick’s theory of knowledge. Synthese, 79(2):305–317, May 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869628.

Balzer:1989:STR

[866] W. Balzer, B. Lauth, and G. Zoubek. A static theory of reference in science. Synthese, 79(3):319–360, June 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869280.

Lai:1989:HWM

[867] Tyrone Lai. How we make discoveries. Synthese, 79(3):361–392, June 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869281.

Woodward:1989:DP

[868] Jim Woodward. Data and phenomena. Synthese, 79(3):393–472, June 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869282.

Lin:1989:MRA

[869] Yi Lin. A multi-relation approach of general systems and tests of applications. Synthese, 79(3):473–488, June 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869283.

Wassermann:1989:TSM

[870] G. D. Wassermann. Theories, systemic models (SYMOs), laws and facts in the sciences. Synthese, 79(3):489–514, June 1989. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869284.

Turney:1989:ACN

[871] Peter Turney. The architecture of complexity: A new blueprint. Synthese, 79(3):515–542, June 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869285.

Gavroglu:1989:SOW

[872] Kostas Gavroğlu. Simplicity and observability: When are particles elementary? Synthese, 79(3):543–557, June 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869286.

Ruger:1989:CMG

[873] Alexander Rüger. Complementarity meets general relativity: A study in ontological commitments and theory unification. Synthese, 79(3):559– 580, June 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869287.

Matthews:1989:CP

[874] Gareth B. Matthews and Thomas A. Blackson. Causes in the Phaedo. Synthese, 79(3):581–591, June 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869288.

Matthews:1989:HPSa

[875] Michael R. Matthews. History, philosophy, and science teaching: A brief review. Synthese, 80(1):1–7, July 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869945.

Siegel:1989:RSC

[876] Harvey Siegel. The rationality of science, critical thinking, and science education. Synthese, 80(1):9–41, July 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869946.

Silverman:1989:TSW

[877] M. P. Silverman. Two sides of wonder: Philosophical keys to the motivation of science learning. Synthese, 80(1):43–61, July 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869947.

Jordan:1989:TSP

[878] Trace Jordan. Themes and schemes: A philosophical approach to interdisciplinary science teaching. Synthese, 80(1):63–79, July 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869948.

Eger:1989:ISP

[879] Martin Eger. The ‘interests’ of science and the problems of education. Synthese, 80(1):81–106, July 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869949.

Benson:1989:MSP

[880] Garth D. Benson. The misrepresentation of science by philosophers and teachers of science. Synthese, 80(1):107–119, July 1989. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869950.

vonGlasersfeld:1989:CCK

[881] Ernst von Glasersfeld. Cognition, construction of knowledge, and teaching. Synthese, 80(1):121–140, July 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869951.

Rowell:1989:PEE

[882] Jack A. Rowell. Piagetian epistemology: Equilibration and the teaching of science. Synthese, 80(1):141–162, July 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869952.

Nersessian:1989:CCS

[883] Nancy J. Nersessian. Conceptual change in science and in science education. Synthese, 80(1):163–183, July 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869953.

Matthews:1989:HPSb

[884] Michael R. Matthews. History, philosophy and science teaching: A bibliography. Synthese, 80(1):185–196, July 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869954.

Anonymous:1989:Ad

[885] Anonymous. Announcement. Synthese, 80(1):199, July 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869955.

Gert:1989:PTC

[886] Bernard Gert. Psychological terms and criteria. Synthese, 80(2):201– 222, August 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869484.

Cunningham:1989:PMM

[887] Suzanne Cunningham. Perception, meaning, and mind. Synthese, 80 (2):223–241, August 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869485.

Tuomela:1989:CAS

[888] Raimo Tuomela. Collective action, supervenience, and constitution. Synthese, 80(2):243–266, August 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869486.

Jacquette:1989:SIT

[889] Dale Jacquette. Searle’s intentionality thesis. Synthese, 80(2):267– 275, August 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869487.

Clark:1989:PIH

[890] Austen Clark. The particulate instantiation of homogeneous pink. Synthese, 80(2):277–304, August 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869488.

Nelson:1989:R

[891] R. J. Nelson. Review. Synthese, 80(2):305–313, August 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869489.

Nelson:1989:CNI

[892] R. J. Nelson. Correction to ‘naturalizing intentions’. Synthese, 80(2): 315–317, August 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869490;   http://link.springer.com/content/pdf/10.1007/BF00869490.
pdf.

Anonymous:1989:PA

[893] Anonymous. Pre-announcement. Synthese, 80(2):319, August 1989. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869491.

Burdick:1989:DI

[894] Howard Burdick. On Davidson and interpretation. Synthese, 80(3):321– 345, September 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869579.

Crumley:1989:TLL

[895] Jack S. Crumley II. Talking lions and lion talk: Davidson on conceptual schemes. Synthese, 80(3):347–371, September 1989. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869580.

Shaw:1989:SUF

[896] J. L. Shaw. ‘Saturated’ and ‘unsaturated’: Frege and the Nyaya. Synthese, 80(3):373–394, September 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869581.

Schwartz:1989:VIR

[897] Stephen P. Schwartz. Vagueness and incoherence: A reply to Burns. Synthese, 80(3):395–406, September 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869582.

Johnson:1989:MFI

[898] Ralph H. Johnson. Massey on fallacy and informal logic: A reply. Synthese, 80(3):407–426, September 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869583.

Hansson:1989:NDS

[899] Sven Ove Hansson. A note on the deontic system DL of Jones and Pörn. Synthese, 80(3):427–428, September 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869584.

Jones:1989:RH

[900] Andrew J. I. Jones and Ingmar Pörn. A rejoinder to hansson. Synthese, 80 (3):429–432, September 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869585.

Carrier:1989:R

[901] L. S. Carrier. Review. Synthese, 80(3):433–446, September 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869586.

Howson:1989:SPB

[902] Colin Howson. Subjective probabilities and betting quotients. Synthese, 81(1):1–8, October 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869341.

Eells:1989:PPS

[903] Ellery Eells. The popcorn problem: Sobel on evidential decision theory and deliberation-probability dynamics. Synthese, 81(1):9–20, October 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869342.

Koons:1989:RAM

[904] Robert C. Koons. A representational account of mutual belief. Synthese, 81(1):21–45, October 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869343.

Halpin:1989:CAC

[905] John F. Halpin. Counterfactual analysis: Can the metalinguistic theory be revitalized? Synthese, 81(1):47–62, October 1989. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869344.

Clendinnen:1989:RUT

[906] F. John Clendinnen. Realism and the underdetermination of theory. Synthese, 81(1):63–90, October 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869345.

Rott:1989:CTC

[907] Hans Rott. Conditionals and theory change: Revisions, expansions, and additions. Synthese, 81(1):91–113, October 1989. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869346.

Fuhrmann:1989:RMT

[908] André Fuhrmann. Reflective modalities and theory change. Synthese, 81 (1):115–134, October 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869347.

Anonymous:1989:Ae

[909] Anonymous. Announcement. Synthese, 81(1):135, October 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869348.

Anonymous:1989:Af

[910] Anonymous. Announcement. Synthese, 81(1):136, October 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869349.

Chihara:1989:GEP

[911] Charles Chihara and Brian Skyrms. Guest editors’ preface. Synthese, 81(2):139, November 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869326.

Wang:1989:TCL

[912] Hao Wang. Tharp and conceptual logic. Synthese, 81(2):141–152, November 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869327.

Chihara:1989:TMM

[913] Charles Chihara. Tharp’s ‘myth and mathematics’. Synthese, 81(2):153– 165, November 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869328.

Tharp:1989:MMC

[914] Leslie Tharp. Myth and mathematics: A conceptualistic philosophy of mathematics I. Synthese, 81(2):167–201, November 1989. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869329.

Skyrms:1989:ITT

[915] Brian Skyrms. Introduction to ‘three theories of metaphysics’. Synthese, 81(2):203–205, November 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869330.

Tharp:1989:TTM

[916] Leslie Tharp. Three theorems of metaphysics. Synthese, 81(2):207– 214, November 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869331.

Anonymous:1989:BLT

[917] Anonymous. Bibliography of Leslie Tharp’s writings. Synthese, 81(2): 215, November 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869332.

Hirsch:1989:NCS

[918] Eli Hirsch. Negativity and complexity: Some logical considerations. Synthese, 81(2):217–241, November 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869333.

Hand:1989:NEO

[919] Michael Hand. A number is the exponent of an operation. Synthese, 81 (2):243–265, November 1989. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869334.

Weil:1989:I

[920] Vivian Weil and Jon Nordby. Introduction. Synthese, 81(3):267–270, December 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869316.

Cartwright:1989:BED

[921] Nancy Cartwright. The Born–Einstein debate: Where application and explanation separate. Synthese, 81(3):271–282, December 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869317.

Denbigh:1989:GPN

[922] K. G. Denbigh and M. L. G. Redhead. Gibbs’ Paradox and non-uniform convergence. Synthese, 81(3):283–312, December 1989. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869318.

Knorr:1989:PEA

[923] Wilbur R. Knorr. The practical element in ancient exact sciences. Synthese, 81(3):313–328, December 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869319.

Shrader-Frechette:1989:ILA

[924] K. S. Shrader-Frechette. Idealized laws, antirealism, and applied science: A case in hydrogeology. Synthese, 81(3):329–352, December 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869320.

Laymon:1989:AIS

[925] Ronald Laymon. Applying idealized scientific theories to engineering. Synthese, 81(3):353–371, December 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869321.

Nordby:1989:BWB

[926] Jon J. Nordby. Bootstrapping while barefoot (crime models vs. theoretical models in the hunt for serial killers). Synthese, 81(3):373–389, December 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869322.

Longino:1989:BEL

[927] Helen E. Longino. Biological effects of low level radiation: Values, doseresponse models, risk estimates. Synthese, 81(3):391–404, December 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869323.

DiNucciPearce:1989:TVS

[928] M. Rosaria Di Nucci Pearce and David Pearce. Technology vs. science: The cognitive fallacy. Synthese, 81(3):405–419, December 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869324.

Anonymous:1989:CP

[929] Anonymous. Call for papers. Synthese, 81(3):421, December 1989. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869325.